### v1

# BATIS: Bootstrapping, Autonomous Testing, and Initialization System for Quantum Dot Devices

Abstract
Semiconductor quantum dot (QD) devices have become central to advancements in spin-based quantum computing. As the complexity of QD devices grows, manual tuning becomes increasingly infeasible, necessitating robust and scalable autotuning solutions. Tuning large arrays of QD qubits depends on efficient choices of automated protocols. Here, we introduce a bootstrapping, autonomous testing, and initialization system (BATIS), an automated framework designed to streamline QD device testing and initialization. BATIS navigates high-dimensional gate voltage spaces, automating essential steps such as leakage testing and gate characterization. The current channel formation protocol follows a novel and scalable approach that requires a single measurement regardless of the number of channels. Demonstrated at on a quad-QD Si/SixGe1-x device, BATIS eliminates the need for deep cryogenic environments during initial device diagnostics, significantly enhancing scalability and reducing setup times. By requiring minimal prior knowledge of the device architecture, BATIS represents a platform-agnostic solution, adaptable to various QD systems, which bridges a critical gap in QD autotuning.
I Introduction
Quantum dot (QD) qubit devices are commonly formed by electrostatically manipulating the potential energy landscape of a two-dimensional (2D) electron or hole gas (EG or HG). This procedure produces interleaved regions of accumulated and depleted charge density that serve as QD qubits, charge sensors, and charge carrier reservoirs [1, 2, 3, 4, 5]. In recent years, the number of QDs in such devices has risen rapidly [6, 7, 8]. In the near term, human intervention will no longer be possible to tune up many dots simultaneously. The field of autotuning has arisen to address this problem.
Modern semiconductor QD qubit devices have two critical features that enhance their performance and simultaneously make autotuning challenging and essential. First, the active region of the device—the QDs, charge sensors, and reservoirs—contains no doping. All charge carriers—electrons or holes—are induced by gate voltages. No current can flow until gate voltages are applied and carriers are accumulated, yet those voltages are unknown a priori. Moreover, the exact voltage configuration changes not only between cooldowns but even between successive initialization steps. Second, QD device chips have multiple current channels that often serve different purposes. Some can host an array of QD qubits, while others provide a home for charge sensors. These channels are not separate until pinch-off voltages, i.e., the gate voltage threshold at which carrier accumulation is eliminated, are determined. Moreover, since the accumulation and depletion of carriers must occur in alternating regions separated by only tens of nanometers to enable the formation of QDs separated by tunnel barriers and proximal charge sensors, these channels must interact.
Here, we describe a bootstrapping, autonomous testing, and initialization system (BATIS) that can automatically test the device and set the gate voltages to ensure that the QD chip is properly accumulated and all required pinch-off voltages are determined. BATIS takes as input the device gate layout and a high-level understanding of the desired operation, including critically where current channels should be continuous and which gates should divide one channel from another. The outputs of BATIS are the gate voltage parameters that make the QD chip functional, with all current channels well delineated and all pinch-off voltages known. Importantly, in this work, the sample is kept at around –about 10 times warmer than other typical devices [9, 10]. Such temperatures are highly desirable for what is known as high-temperature operation [11]. However, higher temperature causes an increased difficulty in thresholding pinchoffs as the coulomb blockade and all measurement features derived therein are less well defined and easier to confuse with noise [12]. BATIS is capable of tuning at these high temperatures, which is important both for scalability and automated testing as, for example, is possible using cryogenic wafer probers [13, 14].
Previous work on autotuning has shown that a bootstrap process is sufficient for further tuning of QD devices. For example, if current channels are known in modulation-doped QD devices, the formation of double QDs with few-electron occupation can be automated [15]. In such devices, autonomous bootstrapping has been demonstrated [16, 17]. If a charge sensor is available, undoped SiMOS devices have been autotuned to achieve a single-electron QD [18]. For nanowires, which by design have a single, known current channel, automation into fully formed QDs has been demonstrated, provided device-specific details are placed into a configuration file [19].
For accumulation-mode, undoped QD devices, previous work has enabled progress in tuning device topology [20, 21] and charge-state calibration [16, 22]. Methods for automated detection of unintentional, spurious QDs have also been proposed [23]. More recently, a framework to achieve orthogonal control of the electrostatic potential landscape in a large 2D array has been put forward [24]. However, autonomous bootstrapping for accumulation mode QD devices has never been attempted, and most tuning algorithms proposed to date assume that steps such as device characterization and the definition of the QD(s) channel(s) have already been completed [25].
The approach for designing BATIS differs from others, typically exercised by the community. All bootstrapping algorithms proposed to date are highly device-specific, with the algorithm’s overall efficiency measured as progress-per-second or some other metric against time used as a figure of merit. However, time-based metrics for device-specific algorithms provide little basis for valid comparison against alternative methods. Such metrics quickly become instrument-bound. Taking large numbers of data points per measurement, each with integration times in the or , quickly dwarfs the simplest point-wise threshold analysis performed on modern multi-core CPUs clocked in the . Users can easily speed up their benchmarks for single-use cases by upgrading to faster instrumentation.
Moreover, the resulting algorithms provide single-use, device-specific applications by assuming substantial knowledge about the tuned state of QD devices. Thus, even if such specialized algorithms perform well on the development device, they require significant redesigning to be deployable on other QD devices.
In contrast, when creating BATIS, we focused on developing a platform-agnostic tool that can be applied to a variety of QD devices. While the procedural flow of BATIS is predetermined, the executable flow is automatically adjusted based on the general device architecture and connectivity to the measurement instruments, both provided within the configuration files. In other words, rather than being programmed to follow a fixed sequence of measurements, BATIS is designed to choose which steps to take and change the expected outcome of each step – depending on the device used. Further adjustments are automatically made as the tuning progresses and the algorithm learns the device.
The device characterization and tuning processes are a repetitive yet essential part of initializing QD-based qubit chips. The majority of autotuning efforts have focused on the latter phases of automatic device tuning, including coarse tuning, charge-state tuning, virtualization, and fine-tuning, which is sufficient for quantum dots in a one-dimensional channel or when the initial step is performed by hand [25]. In modern chips, the scale of the tune-up will soon preclude hand-tuning as fully 2D structures become ubiquitous. BATIS addresses the challenges that are about to become limiting in this widespread class of quantum semiconductor chips.
II BATIS Design
Figure 1(a) shows a gate-defined QD device that illustrates the critical features of accumulation-mode, undoped devices: multiple channels (three white arrows), regions for forming quantum dots (four in the qubit channel, below, and two single-QD charge sensors, above), and reservoir gates to connect these regions to ohmic contacts for current flow. This device exemplifies a commonly used fabrication strategy known as overlapping gates [27, 28].
The screening gates, shown in yellow, are the closest layer of gates to the top gate oxide in the three-layer gate design and are used to isolate each of the channels from each other and from other regions in the chip. The second layer (shown in blue and red) includes the reservoir and plunger gates. The reservoirs enable the flow of electrons from the ohmic contacts toward the center of the device where QDs will form in the electrostatic potential profile controlled by each plunger gate. The final layer consists of barrier gates (grey), which control the tunnel couplings between QDs—both qubit QDs and charge-sensing QDs—and between the outermost QDs and reservoirs. By carefully adjusting the voltages applied to each of the gates, the quantum dot qubit chip with charge sensors is made functional [2, 3, 4, 5].
BATIS is designed to automate the task of testing and characterizing all gates and forming all necessary 1D current channels for arbitrary-length linear QD arrays. The procedural flow of the single button press algorithm is depicted in Fig. 1(b).
BATIS takes as input two configuration files describing the device design and certain generic characteristics of the device and instruments. Based on that information, it performs full diagnostics and characterization of each gate and determines the proper configuration for the consecutive stage of the tuning processes. Each of the six stages in BATIS involves at least one device measurement followed by data analysis to test the functionality of a selected subset of gates, shown in Fig. 1(b) by gray rounded rectangles. The optional illumination stage is executed on an as-needed basis. Illumination involves exposing the QD device to 780 light under carefully chosen gate voltages, to modify its state. Such exposure shifts the gate voltage response by modifying the amount (and if desired, the sign) of charge trapped at the semiconductor-oxide interface [26].
The algorithm follows a conditional logic visually represented with arrows of different colors and textures in Fig. 1(b). The solid gray arrows indicate the procedural step in the sequence. Passing a test for a current stage, marked with yellow diamonds in Fig. 1(b), initiates a successive measurement and assessment stage, as indicated with solid green arrows. Failing a test can result in repeated measurements, device illumination, or tuning termination, depending on the stage. The measurement schemes, together with idealized and real-world data associated with each stage, are shown in Fig. 2.
Once completed, BATIS produces a device diagnostics file that stores all information about each gate learned during tuning. It also pre-configures the QD device for further use. The details of the flow of the algorithm are discussed in the following sections.
II.1 Initialization and First Leakage Test
The two configuration files used as input for BATIS include descriptions of all gates, desirable current channels, and the QD device connections. The files include also details about the software configuration necessary to execute BATIS and a list of variables that need to be considered at each stage. These variables consist of instrumentation limits to ensure that the algorithm does not search in ranges not supported by the instruments and device limits to prevent the device from breaking during tune-up. Details included in configuration files allow the linking of abstract concepts about the device components with the language used in BATIS.
The functionality of gates is determined automatically from the device architecture that the end user inputs, building a gate-lookup table. The files also contain a wire map and instrument information detailing the measurement system setup such that BATIS can automatically read and execute all specialized measurements, connecting the gate-lookup table with the measurement instruments. This allows BATIS to remain device- and instrument-agnostic.
BATIS starts by checking the device for leakage to the fridge ground through the oxide. The first leakage measurement happens before the electron accumulation during the leakage test stage, see Fig. 1(b). The resistances of each connection are measured with respect to each other and compiled into a leakage matrix, as shown in the second and third row of Fig. 2(a). The desired result is a high resistance between all the connections, an entirely yellow matrix.
Defects during the fabrication process can result in leakage. Leakage manifests as off-colored cells in the matrix, as shown in the right panel in the second row and in the third row of Fig. 2(a). Should leakage be detected, BATIS terminates and flags the device as unusable.
II.2 Global Accumulation
For accumulation mode QD devices no 2DEG is present at zero applied gate voltage. Applying a non-zero bias lowers the conduction band minimum below the Fermi level, accumulating a 2DEG in the quantum well near the Si/SixGe1-x interface. The goal of the global accumulation stage, see Fig. 1(b), is to determine the characteristic voltage where enough of the 2DEG is accumulated that a significant amount of current begins to flow between the ohmic contacts, effectively turning the device on.
Idealized successful and failing measurements for this stage are visualized in the left and right panels in the middle row of Fig. 2(b), respectively. The turn-on process in real QD devices is not always straightforward, as even a good device might not turn on properly when first tested. If the algorithm encounters a problem within global turn-on, it has a recovery pathway through optional illumination, see Fig. 1(b).
Illumination is initiated when the device fails to turn on or the turn-on voltage is much too high to be usable. After illumination, the global accumulation stage is re-initiated. If the preference is not to use illumination, BATIS allows skipping this step and going straight to repeating the global turn-on stage. Whether or not illumination should be used is an option that is pre-defined in the configuration file.
II.3 Accumulated Leakage Test
Once the 2DEG is accumulated, the second, accumulated leakage test is initiated. This step proceeds analogously to the first leakage check, except that it does not include the ohmic connections. This is because this leakage check is taken when the device is accumulated, and the ohmics are intentionally shorted to each other. If this stage fails, it indicates that a gate is shorted to the 2DEG flowing beneath it, which is an unrecoverable fabrication defect. Should that occur, BATIS terminates.
II.4 Screening and Reservoir Characterization
The turn-on characteristic voltage is taken globally, meaning it does not provide gate-specific information. To learn that information, BATIS performs a series of individual gate characterizations. The first of these is performed at the screening and reservoir characterization stage in Fig. 1(b).
With the 2DEG accumulated and the device turned on, BATIS depletes the voltages on individual gates until the current on the corresponding channel is pinched off, as displayed in the right panel in the second row in Fig. 2(c). BATIS uses these pinch-off curves to determine where to place reservoir gates for proper 2DEG accumulation and establish bounds to operate the screening gates.
The typical failure modes for this stage include too much depletion and too much accumulation. If the channel below a gate is extraordinarily depleted, no voltage can be applied to the gate to accumulate it. However, since the screening and reservoir characterization stage is only initialized if all current channels turn on properly at the global accumulation stage, the inability to accumulate implies either an issue with the measurement instrument or hysteresis with the gate’s action upon the 2DEG. Similarly, pinching off of all current channels is also confirmed at the global accumulation stage, and thus, it is implausible to observe too much accumulation at this stage. To test whether the issue is caused by the instrument, the algorithm executes a repeated measurement. If the error is observed a second time, it indicates an issue with the gate and the algorithm terminates.
II.5 Forming 1D Current Channels
Selectively removing 2DEG from certain regions of the device is done to form the 1D current channels. It requires depleting the screening gates and accumulating the plungers and barriers (often collectively called finger gates) such that the current flows between the screening gates along a desirable path, see the left panel in Fig. 2(c).
Determining voltages where 1D channels are properly formed is typically done by sweeping two screening gates simultaneously in a 2D sweep while holding the finger gates constant. The formation of the channel manifests as a triangular-sloped region (the so-called triangle plots). If the finger gates are not accumulated enough, this region will not form. Thus, multiple scans with incremented finger gate voltages—a three-dimensional dataset—are often required before the region is observed for all required channels.
BATIS employs a different approach to simultaneously accumulate multiple 1D channels using only a single two-dimensional dataset. Instead of sweeping two screening gates, BATIS takes a 2D voltage scan while sweeping one screening gate against all of the finger gates between that gate and the other screening gate used to form the 1D channel (typically the central screening gate). Since BATIS already learned the necessary configuration for the channel isolation with the central screening gate, this information is now utilized to configure the remaining screening gates. The 1D channel formation technique implemented in BATIS requires a single measurement for each pair of gates, which makes it much more efficient than the standard approaches.
The plot in the right panel in the second row Fig. 2(d) exhibits three characteristic regions: a no-current region in the bottom left, a screening-gate-dominated region in the bottom right, and a finger-gate-dominated region in the top left. BATIS automatically selects an operating point in the top left part of the plot, near the boundary between the current and the no-current regions. The automatic analysis of the regions and their respective boundaries is enabled by fitting a synthetic model to the experimental data, similar to the presentation in [29]. Additional criteria based on the local properties, e.g., estimated resistance and derivatives of the measurement, are used to optimize the selection of the final point among numerous candidates in the operating region.
If the finger-gate-dominated region does not form, BATIS illuminates the device with the finger gates at a voltage 400 more negative than the most recent illumination. After such an illumination, the entire device response can change, because of cross-talk and the interdependence of the current paths. Thus, BATIS resumes autotuning from the global accumulation step.
II.6 Finger Gate Characterization
Once isolated current channels are formed, finger gate pinch-offs can be measured. This is done by assuming that only the nearest neighbors can affect the accumulation under a gate via crosstalk when pinching off the channels. Ideally, as the voltage on a given finger gate is lowered, the pinch-off curves should exhibit a smooth, sigmoidal trend, eventually cutting off the current in the respective channels. An idealized representation of such a curve is visualized in the left panel in the second row in Fig. 2(e).
The bottom panel in the third row of Fig. 2(e) shows a typical curve for these measurements. As the voltage is reduced, the curve begins to exhibit the characteristic Coulomb blockade oscillations typical of QD devices [12]. In principle, such oscillations could complicate any fit-based analysis. Instead, BATIS uses thresholding and slope detection to determine the pinch-off voltage on a point-by-point basis, going in the direction of increasing voltage (the reverse of how the signal was measured) to evade the small deviations of background noise, whilst capturing the first significant turn-on.
III Experimental Demonstration
We now demonstrate BATIS on a 1D quadruple Si/SixGe1-x QD device, nominally identical to the one shown in Fig. 1(a). The device’s schematic consisting of 24 gates is shown in Fig. 3(a). The first layer of gates is made up of four screening gates, Si for , shown in Fig. 3(a) in yellow. The next layer consists of five reservoirs, Ri for , and six plunger gates, Pi for . The reservoirs are shown in Fig. 3(a) in blue, while the plungers are shown in red. The final layer consists of nine barrier gates, Bi for , shown in Fig. 3(a) in gray.
The goal of BATIS is to form three current channels in the quad-QD device, as indicated with white arrows in Fig. 3(a). BATIS brings a grounded but not characterized device from Fig. 1(a) to a fully operational device in Fig. 1(d). In what follows, we explain how BATIS performs the quad-QD device bootstrapping, testing, and characterizations.
First Leakage Test. BATIS begins by performing a device leakage test. For the quad-QD device used in our demonstration, the sample mount holding it contains connections, of which are used for instrument channels to control individual gates and ohmics and an additional wired channels are unused for this device. Thus, the size of the leakage matrix is to examine all possible leakage states. BATIS first checks the diagonal elements of the leakage matrix. If the resistance of every connection exceeds 25, BATIS declares that there is no leakage in the device.
If any of the diagonal elements falls below the 25 threshold, the device is declared leaking. Depending on where the current leak occurs, it is sometimes possible to recover a leaking device. Thus, rather than terminating, BATIS continues to fill in the leakage matrix for diagnostic purposes. Due to the symmetry of the measurements, the leakage matrix is also symmetric and only the lower triangular part of the matrix needs to be completed. Moreover, passing the threshold test by a diagonal component automatically guarantees that the entire column corresponding to that component also passes the test. Thus, only the columns corresponding to the elements below 25 need to be determined. BATIS continues to fill the leakage matrix one column at a time. In the worst case, a total of resistance measurements are necessary to complete the leakage matrix.
The leakage test needs to be executed only once per cool-down if under fully automatic control. This is because BATIS has built-in safety checks during voltage choice selection to avoid exceeding safe voltage breakdowns between neighboring gates for all measurements.
Global Accumulation. Once no leakage is confirmed, BATIS performs the global turn-on stage. In our demonstration, the algorithm executed three rounds of illumination before the observed global turn-on values were determined to be satisfactory, occurring between and . During the first two illuminations, all gates were grounded to reset the device after cooldown and verify that no instruments were erratic. After that, for the third illumination, BATIS chose to apply voltages to all screening and reservoir gates. The voltages required to achieve turn-ons at around , calculated following the protocol in Ref. [26], were set to for gates associated with channel I1 and I2, and for gates associated with channel I3. The central screening gate, S2, was left grounded during the illumination.
The fourth iteration of global turn-on measurement was successful. The summary of all four runs is presented in Table 1. The resulting configuration of the QD device is shown in the left panel in Fig. 1(c), where the pinch-offs on all the screening gates, excluding S2, and all of the reservoir gates are colored based on the learned gate voltage.
| Channel | Run 1 | Run 2 | Run 3 | Run 4 | |
| characteristics | |||||
| I1 | VP () | ||||
| Vmax () | |||||
| I2 | VP () | ||||
| Vmax () | |||||
| I3 | VP () | ||||
| Vmax () | |||||
Accumulated Leakage Test. With the global accumulation confirmed, BATIS proceeds to the accumulated leakage test. A second leakage matrix is determined, following the same strategy as during the first leakage test. Like previously, BATIS finds that all resistances exceed 25 . Since the ohmic connections are excluded, the resulting matrix has five fewer indexes. As none of the connections leaked, only measurements following the matrix diagonal were taken.
Screening and Reservoir Characterization. The next stage in BATIS involves the characterization of the screening and reservoir gates. The results of these characterizations for all tested gates are presented in Fig. 3(b). Unlike the other screening and reservoir gates, S2 needs to be measured simultaneously on channels I1, I2, and I3 since it borders all of the current channels. Because of this, the data for the S2 pinch-off can be found in Fig. 3(c) as it needs to be organized by channel.
| Gate | VP () | VI () | VO () |
| S2 | |||
| S1 | |||
| R1 | |||
| R2 | |||
| S3 | |||
| R3 | |||
| R4 | |||
| S4 | |||
| R4 | |||
| R5 |
The voltage values are quite consistent between the sets of reservoir and screening gates intended to form the desired current channels, a feature from the illumination from earlier [26]. We note that for gate S1, BATIS acquired signal over , i.e., twice as long as for all remaining gates. This is likely due to poor fit scoring using the reduced voltage scan range, which caused BATIS to perform additional measurements with an increased sweep range to be more confident in the derived pinch-off value. Since the long tail of the signal has no features, the signal shown in Fig. 3(b) is truncated to be consistent with the remaining measurements for visualization purposes.
BATIS sets the reservoir gates two standard deviations above the mean of the phenomenological fit it performed on the signal’s gradient. This is to ensure that reservoirs will consistently provide electron flow through the channel as we do not want the cross-talk effects to cause drastic changes in the electron accumulation. At the same time, if the reservoirs are too accumulated, they can form spurious channels behind the regions of interest. To resolve this, the screening gates are set below their pinch-offs in Fig. 3(b) to prevent current flow through spurious channels 111The derivative of the screening gate pinch-off was fit to a function like . The non-central screening gates were set to be before the mean of the derivative..
The summary of the pinch-off voltages and the operating point selected by BATIS is presented in Table 2. The resulting configuration, where the device is in the correct operating point to perform channel accumulation, is visualized in the central plot in Fig. 1(c).
Forming 1D Current Channels. With all screening and reservoir gates characterized, BATIS proceeds to form the 1D current channels. The quad-QD device requires three current channels to be formed, as indicated with the white arrows in Fig. 3(a). BATIS identified three different operating points, one for each 1D current channel calculated from the simultaneous channel accumulation in Fig. 3(c). The summary of the screening and finger gates configuration for each operating point is presented in Table 3, and a collapsed zero cross-talk visualization is pictured in the right panel in Fig. 1(c). The three points are consolidated into one where the voltage color for each gate is the VO indexed by the channel the gate belongs to.
| Gate | V () | V () | V () |
| S1 | |||
| S3 | |||
| S4 | |||
As we discuss in Sec.II.5, the algorithm for 1D current channel formation implemented in BATIS does not follow the typical, triangle-plot-based strategy. Rather, it sweeps each of the three outer screening gates, i.e., S1, S3, and S4, against all finger gates relevant to the given 1D current channel. The traditional triangle plots [31, 29] are orthogonal to the measured planes shown in Fig. 3(c) and would have been measurements with the S2 axis instead. In other words, the triangle plots and the plots utilized in BATIS contain similar information on different plane cuts through the same space.
The accumulation plots in Fig. 3(c) reveal a dip in the current exiting when the current starts to flow in , shown by the dark green shape imposed on top of the signal in . This is due to a coupling between those two channels resulting from a shared ohmic connection to source the currents. Such a dip in the charge sensor readout is often called a shadow. Shadows increase the difficulty of detecting and forming isolated current channels. The accumulation plot for I1 is much less correlated with the current flow in the two charge sensor channels, but there is still some cross-talk. This is due to separation with the S2 gate which divides the 2DEG between the upper readout channels and the lower dot channel.
Finger Gate Characterization. The last stage of BATIS involves finger gate characterization. This characterization is done individually for all of the finger gates on the QD device. The voltage range for each measurement is always the same and predefined in the configuration file. However, due to the Coulomb blockade oscillations, the automated search for the starting voltage for some gates is more challenging than for others. This can be seen in the raw data pictured in Fig. 3(d).
To ensure that the right point is consistently found, BATIS combines thresholding and slope extraction for each plot to ensure that the pinch-off it finds is robust to noise and spurious fluctuations that do not exceed the threshold. The resulting pinch-off values of all of these varying curves for all finger gates are presented in Table 4. This configuration is used in the final colorized SEM for the ending state of the algorithm shown in Fig. 1(d).
| Gate | VP () | Gate | VP () | Gate | VP () |
| B1 | B6 | B8 | |||
| P1 | P5 | P6 | |||
| B2 | B7 | B9 | |||
| P2 | |||||
| B3 | |||||
| P3 | |||||
| B4 | |||||
| P4 | |||||
| B5 | |||||
IV Summary and outlook
In this paper, we proposed and demonstrated BATIS: a fully automatic procedure for bootstrapping, automated testing, and characterization of accumulation mode, undoped QD devices. The prior knowledge about the QD device required to initialize BATIS is limited to the heterostructure, the device geometry, and the device limits. BATIS navigates through complex high-dimensional gate space to find the final gate configuration in which all desirable current channels are formed following simple physical principles.
The goal of BATIS is to test the device’s functionality to confirm that all gates can properly pinch off and that the current channels can be formed. Testing and characterizing an unknown QD device is essential for identifying suitable global device states that enable the observation of Coulomb blockade. Importantly, the single button press testing and initialization is demonstrated at , eliminating the need for a full cooldown just to learn that a device is not functional. This result is particularly important as we envision a hot large-scale spin-based quantum processor. These systems will require qubit initialization, relying on a complete and accurate device characterization. One could imagine using BATIS to probe the attributes of many devices spread across a wafer or provide device diagnostics on a partition of a large qubit processor without going to temperatures.
The flexible structure allows BATIS to be scalable to characterize and initialize any accumulation mode linear array of QD and easily adaptable to other QD platforms. In particular, BATIS could straightforwardly mimic the behavior of algorithms proposed for nanowires and GaAs/AlxGa1-xAs by simply invoking skip logic on the steps not necessarily supported for these devices. However, it is quite unlikely that any previously proposed algorithms could be easily adapted to the Si/SixGe1-x quad-QD device considered in this work.
Previous studies have demonstrated a device-agnostic algorithm for certain aspects of QD device tuning, such as setting topology [32, 17] and charge state classification [33, 22]. However, all these algorithms presuppose that device characterization and foundational tuning steps—determining gate functionality, defining the QD channel(s), or locating regions of interest for the coarse tuning regime—have already been performed. BATIS fills this gap, making device characterization easily within reach for linear QD array systems and paving the way toward fully automatic control. When linked together, all proposed subroutines can form an autonomous system for controlling arbitrary QD devices up to any desired level of characterization, which will accommodate the growth of QD qubit systems on the path to a scalable quantum computer on this platform.
Acknowledgements.
We acknowledge Danielle Middlebrooks for helping with the automation. We acknowledge HRL Laboratories, LLC for support and L.F. Edge for providing one of the Si/SixGe1-x heterostructures used in this work. This work supported in part by ARO grant nos. W911NF-23-1-0110 and W911NF-17-1-0274. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright noted herein. Any mention of commercial products is for information only; it does not imply recommendation or endorsement by NIST.References
- Burkard et al. [2023] G. Burkard, T. D. Ladd, A. Pan, J. M. Nichol, and J. R. Petta, Semiconductor spin qubits, Reviews of Modern Physics 95, 025003 (2023).
- van der Wiel et al. [2002] W. G. van der Wiel, S. De Franceschi, J. M. Elzerman, T. Fujisawa, S. Tarucha, and L. P. Kouwenhoven, Electron transport through double quantum dots, Rev. Mod. Phys. 75, 1 (2002).
- Loss and DiVincenzo [1998] D. Loss and D. P. DiVincenzo, Quantum computation with quantum dots, Phys. Rev. A 57, 120 (1998).
- Hanson et al. [2007] R. Hanson, L. P. Kouwenhoven, J. R. Petta, S. Tarucha, and L. M. K. Vandersypen, Spins in few-electron quantum dots, Rev. Mod. Phys. 79, 1217 (2007).
- Zwanenburg et al. [2013] F. A. Zwanenburg, A. S. Dzurak, A. Morello, M. Y. Simmons, L. C. L. Hollenberg, G. Klimeck, S. Rogge, S. N. Coppersmith, and M. A. Eriksson, Silicon quantum electronics, Rev. Mod. Phys. 85, 961 (2013).
- Ha et al. [2022] W. Ha, S. D. Ha, M. D. Choi, Y. Tang, A. E. Schmitz, M. P. Levendorf, K. Lee, J. M. Chappell, T. S. Adams, D. R. Hulbert, E. Acuna, R. S. Noah, J. W. Matten, M. P. Jura, J. A. Wright, M. T. Rakher, and M. G. Borselli, A flexible design platform for si/sige exchange-only qubits with low disorder, Nano Lett. 22, 1443 (2022), pMID: 34806894.
- Wang et al. [2024] C.-A. Wang, V. John, H. Tidjani, C. X. Yu, A. S. Ivlev, C. Déprez, F. van Riggelen-Doelman, B. D. Woods, N. W. Hendrickx, W. I. L. Lawrie, L. E. A. Stehouwer, S. D. Oosterhout, A. Sammak, M. Friesen, G. Scappucci, S. L. de Snoo, M. Rimbach-Russ, F. Borsoi, and M. Veldhorst, Operating semiconductor quantum processors with hopping spins, Science 385, 447 (2024).
- George et al. [2024] H. C. George, M. T. Mądzik, E. M. Henry, A. J. Wagner, M. M. Islam, F. Borjans, E. J. Connors, J. Corrigan, M. Curry, M. K. Harper, D. Keith, L. Lampert, F. Luthi, F. A. Mohiyaddin, S. Murcia, R. Nair, R. Nahm, A. Nethwewala, S. Neyens, R. D. Raharjo, C. Rogan, R. Savytskyy, T. F. Watson, J. Ziegler, O. K. Zietz, R. Pillarisetty, N. C. Bishop, S. A. Bojarski, J. Roberts, and J. S. Clarke, 12-spin-qubit arrays fabricated on a 300 mm semiconductor manufacturing line (2024), arXiv:2410.16583 [cond-mat.mes-hall] .
- Borsoi et al. [2024] F. Borsoi, N. W. Hendrickx, V. John, M. Meyer, S. Motz, F. Van Riggelen, A. Sammak, S. L. De Snoo, G. Scappucci, and M. Veldhorst, Shared control of a 16 semiconductor quantum dot crossbar array, Nature Nanotechnology 19, 21 (2024).
- Park et al. [2024] S. Park, J. Benson, J. Corrigan, J. P. Dodson, S. N. Coppersmith, M. Friesen, and M. A. Eriksson, Single-shot latched readout of a quantum dot qubit using barrier gate pulsing (2024), arXiv:2408.15380 [cond-mat].
- Vandersypen et al. [2017] L. M. K. Vandersypen, H. Bluhm, J. S. Clarke, A. S. Dzurak, R. Ishihara, A. Morello, D. J. Reilly, L. R. Schreiber, and M. Veldhorst, Interfacing spin qubits in quantum dots and donors—hot, dense, and coherent, npj Quantum Inf. 3, 34 (2017).
- Meirav et al. [1990] U. Meirav, M. A. Kastner, and S. J. Wind, Single-electron charging and periodic conductance resonances in gaas nanostructures, Phys. Rev. Lett. 65, 771 (1990).
- de Kruijf et al. [2023] M. de Kruijf, S. Geyer, T. Berger, M. Mergenthaler, F. Braakman, R. J. Warburton, and A. V. Kuhlmann, A compact and versatile cryogenic probe station for quantum device testing, Review of Scientific Instruments 94, 054707 (2023).
- Neyens et al. [2024] S. Neyens, O. K. Zietz, T. F. Watson, F. Luthi, A. Nethwewala, H. C. George, E. Henry, M. Islam, A. J. Wagner, F. Borjans, E. J. Connors, J. Corrigan, M. J. Curry, D. Keith, R. Kotlyar, L. F. Lampert, M. T. Mądzik, K. Millard, F. A. Mohiyaddin, S. Pellerano, R. Pillarisetty, M. Ramsey, R. Savytskyy, S. Schaal, G. Zheng, J. Ziegler, N. C. Bishop, S. Bojarski, J. Roberts, and J. S. Clarke, Probing single electrons across 300-mm spin qubit wafers, Nature 629, 80 (2024).
- Baart et al. [2016] T. A. Baart, P. T. Eendebak, C. Reichl, W. Wegscheider, and L. M. K. Vandersypen, Computer-automated tuning of semiconductor double quantum dots into the single-electron regime, Appl. Phys. Lett. 108, 213104 (2016).
- Darulová et al. [2020] J. Darulová, S. J. Pauka, N. Wiebe, K. W. Chan, G. C. Gardener, M. J. Manfra, M. C. Cassidy, and M. Troyer, Autonomous tuning and charge-state detection of gate-defined quantum dots, Phys. Rev. Appl. 13, 054005 (2020).
- Zubchenko et al. [2024] A. Zubchenko, D. Middlebrooks, T. Rasmussen, L. Lausen, F. Kuemmeth, A. Chatterjee, and J. P. Zwolak, Autonomous bootstrapping of quantum dot devices, arXiv:2407.20061 (2024).
- Lapointe-Major et al. [2020] M. Lapointe-Major, O. Germain, J. Camirand Lemyre, D. Lachance-Quirion, S. Rochette, F. Camirand Lemyre, and M. Pioro-Ladrière, Algorithm for automated tuning of a quantum dot into the single-electron regime, Phys. Rev. B 102, 085301 (2020).
- Severin et al. [2024] B. Severin, D. T. Lennon, L. C. Camenzind, F. Vigneau, F. Fedele, D. Jirovec, A. Ballabio, D. Chrastina, G. Isella, M. De Kruijf, M. J. Carballido, S. Svab, A. V. Kuhlmann, S. Geyer, F. N. M. Froning, H. Moon, M. A. Osborne, D. Sejdinovic, G. Katsaros, D. M. Zumbühl, G. A. D. Briggs, and N. Ares, Cross-architecture tuning of silicon and SiGe-based quantum devices using machine learning, Sci. Rep. 14, 17281 (2024).
- Zwolak et al. [2020] J. P. Zwolak, T. McJunkin, S. S. Kalantre, J. Dodson, E. R. MacQuarrie, D. Savage, M. Lagally, S. Coppersmith, M. A. Eriksson, and J. M. Taylor, Autotuning of double-dot devices in situ with machine learning, Phys. Rev. Appl. 13, 034075 (2020).
- Zwolak et al. [2021] J. P. Zwolak, T. McJunkin, S. S. Kalantre, S. F. Neyens, E. R. MacQuarrie, M. A. Eriksson, and J. M. Taylor, Ray-based framework for state identification in quantum dot devices, PRX Quantum 2, 020335 (2021).
- Ziegler et al. [2023a] J. Ziegler, F. Luthi, M. Ramsey, F. Borjans, G. Zheng, and J. P. Zwolak, Tuning arrays with rays: Physics-informed tuning of quantum dot charge states, Phys. Rev. Appl. 20, 034067 (2023a).
- Ziegler et al. [2023b] J. Ziegler, F. Luthi, M. Ramsey, F. Borjans, G. Zheng, and J. P. Zwolak, Automated extraction of capacitive coupling for quantum dot systems, Phys. Rev. Appl. 19, 054077 (2023b).
- Rao et al. [2024] A. S. Rao, D. Buterakos, B. van Straaten, V. John, C. X. Yu, S. D. Oosterhout, L. Stehouwer, G. Scappucci, M. Veldhorst, F. Borsoi, and J. P. Zwolak, MAViS: Modular Autonomous Virtualization System for Two-Dimensional Semiconductor Quantum Dot Arrays, arXiv:2411.12516 (2024).
- Zwolak and Taylor [2023] J. P. Zwolak and J. M. Taylor, Colloquium: Advances in automation of quantum dot devices control, Rev. Mod. Phys. 95, 011006 (2023).
- Wolfe et al. [2024] M. A. Wolfe, B. X. Coe, J. S. Edwards, T. J. Kovach, T. McJunkin, B. Harpt, D. Savage, M. G. Lagally, R. McDermott, M. Friesen, S. Kolkowitz, and M. A. Eriksson, Control of threshold voltages in Si/Si0.7 Ge0.3 quantum devices via optical illumination, Physical Review Applied 22, 034044 (2024).
- Zajac et al. [2016] D. M. Zajac, T. M. Hazard, X. Mi, E. Nielsen, and J. R. Petta, Scalable gate architecture for a one-dimensional array of semiconductor spin qubits, Phys. Rev. Appl. 6, 054013 (2016).
- Dodson et al. [2020] J. P. Dodson, N. Holman, B. Thorgrimsson, S. F. Neyens, E. R. MacQuarrie, T. McJunkin, R. H. Foote, L. F. Edge, S. N. Coppersmith, and M. A. Eriksson, Fabrication process and failure analysis for robust quantum dots in silicon, Nanotechnology 31, 505001 (2020).
- Schug et al. [2024] D. Schug, T. J. Kovach, M. A. Wolfe, J. Benson, S. Park, J. P. Dodson, J. Corrigan, M. A. Eriksson, and J. P. Zwolak, Automation of Quantum Dot Measurement Analysis via Explainable Machine Learning (2024), version Number: 4.
- Note [2] The derivative of the screening gate pinch-off was fit to a function like . The non-central screening gates were set to be before the mean of the derivative.
- McJunkin [2021] T. W. McJunkin, Heterostructure Modifications, Fabrication Improvements, and Measurements Automation of Si/SiGe Quantum Dots for Quantum Computation, Ph.D. thesis, The University of Wisconsin-Madison, Madison, WI, USA (2021).
- Ziegler et al. [2022] J. Ziegler, T. McJunkin, E. S. Joseph, S. S. Kalantre, B. Harpt, D. E. Savage, M. G. Lagally, M. A. Eriksson, J. M. Taylor, and J. P. Zwolak, Toward robust autotuning of noisy quantum dot devices, Phys. Rev. Appl. 17, 024069 (2022).
- Durrer et al. [2020] R. Durrer, B. Kratochwil, J. Koski, A. Landig, C. Reichl, W. Wegscheider, T. Ihn, and E. Greplova, Automated tuning of double quantum dots into specific charge states using neural networks, Phys. Rev. Appl. 13, 054019 (2020).

### v2

@@
 Tuning large arrays of QD qubits depends on efficient choices of automated protocols.
-Here, we introduce a bootstrapping, autonomous testing, and initialization system (BATIS), an automated framework designed to streamline QD device testing and initialization.
+Here, we introduce a bootstrapping, autonomous testing, and initialization system (BATIS) designed to streamline QD device evaluation and calibration.
 BATIS navigates high-dimensional gate voltage spaces, automating essential steps such as leakage testing and gate characterization.
-The current channel formation protocol follows a novel and scalable approach that requires a single measurement regardless of the number of channels.
+For forming the current channels, BATIS follows a non-standard approach that requires a single measurement regardless of the number of channels.
 Demonstrated at on a quad-QD Si/SixGe1-x device, BATIS eliminates the need for deep cryogenic environments during initial device diagnostics, significantly enhancing scalability and reducing setup times.
-By requiring minimal prior knowledge of the device architecture, BATIS represents a platform-agnostic solution, adaptable to various QD systems, which bridges a critical gap in QD autotuning.
+By requiring only minimal prior knowledge of the device architecture, BATIS represents a platform-agnostic solution, adaptable to various QD systems, which bridges a critical gap in QD autotuning.
 I Introduction
@@
 In recent years, the number of QDs in such devices has risen rapidly [6, 7, 8].
-In the near term, human intervention will no longer be possible to tune up many dots simultaneously.
-The field of autotuning has arisen to address this problem.
-Modern semiconductor QD qubit devices have two critical features that enhance their performance and simultaneously make autotuning challenging and essential.
+In the near term, it will no longer be possible to tune up many dots simultaneously using human-based techniques.
+The field of autotuning has arisen to address this challenge.
+Modern semiconductor QD qubit devices have two critical features that enhance their performance, while simultaneously making autotuning challenging and essential.
 First, the active region of the device—the QDs, charge sensors, and reservoirs—contains no doping.
 All charge carriers—electrons or holes—are induced by gate voltages.
-No current can flow until gate voltages are applied and carriers are accumulated, yet those voltages are unknown a priori.
-Moreover, the exact voltage configuration changes not only between cooldowns but even between successive initialization steps.
+No current can flow until gate voltages are applied and carriers are accumulated, yet turn-on voltages are unknown a priori.
+Furthermore, because of device drift and hysteresis, the exact voltage configuration changes not only between cooldowns but even between successive initialization steps.
 Second, QD device chips have multiple current channels that often serve different purposes.
 Some can host an array of QD qubits, while others provide a home for charge sensors.
-These channels are not separate until pinch-off voltages, i.e., the gate voltage threshold at which carrier accumulation is eliminated, are determined.
-Moreover, since the accumulation and depletion of carriers must occur in alternating regions separated by only tens of nanometers to enable the formation of QDs separated by tunnel barriers and proximal charge sensors, these channels must interact.
-Here, we describe a bootstrapping, autonomous testing, and initialization system (BATIS) that can automatically test the device and set the gate voltages to ensure that the QD chip is properly accumulated and all required pinch-off voltages are determined.
-BATIS takes as input the device gate layout and a high-level understanding of the desired operation, including critically where current channels should be continuous and which gates should divide one channel from another.
-The outputs of BATIS are the gate voltage parameters that make the QD chip functional, with all current channels well delineated and all pinch-off voltages known.
-Importantly, in this work, the sample is kept at around –about 10 times warmer than other typical devices [9, 10].
-Such temperatures are highly desirable for what is known as high-temperature operation [11].
-However, higher temperature causes an increased difficulty in thresholding pinchoffs as the coulomb blockade and all measurement features derived therein are less well defined and easier to confuse with noise [12].
-BATIS is capable of tuning at these high temperatures, which is important both for scalability and automated testing as, for example, is possible using cryogenic wafer probers [13, 14].
-Previous work on autotuning has shown that a bootstrap process is sufficient for further tuning of QD devices.
-For example, if current channels are known in modulation-doped QD devices, the formation of double QDs with few-electron occupation can be automated [15].
-In such devices, autonomous bootstrapping has been demonstrated [16, 17].
-If a charge sensor is available, undoped SiMOS devices have been autotuned to achieve a single-electron QD [18].
-For nanowires, which by design have a single, known current channel, automation into fully formed QDs has been demonstrated, provided device-specific details are placed into a configuration file [19].
-For accumulation-mode, undoped QD devices, previous work has enabled progress in tuning device topology [20, 21] and charge-state calibration [16, 22].
-Methods for automated detection of unintentional, spurious QDs have also been proposed [23].
-More recently, a framework to achieve orthogonal control of the electrostatic potential landscape in a large 2D array has been put forward [24].
-However, autonomous bootstrapping for accumulation mode QD devices has never been attempted, and most tuning algorithms proposed to date assume that steps such as device characterization and the definition of the QD(s) channel(s) have already been completed [25].
-The approach for designing BATIS differs from others, typically exercised by the community.
-All bootstrapping algorithms proposed to date are highly device-specific, with the algorithm’s overall efficiency measured as progress-per-second or some other metric against time used as a figure of merit.
-However, time-based metrics for device-specific algorithms provide little basis for valid comparison against alternative methods.
-Such metrics quickly become instrument-bound.
-Taking large numbers of data points per measurement, each with integration times in the or , quickly dwarfs the simplest point-wise threshold analysis performed on modern multi-core CPUs clocked in the .
-Users can easily speed up their benchmarks for single-use cases by upgrading to faster instrumentation.
-Moreover, the resulting algorithms provide single-use, device-specific applications by assuming substantial knowledge about the tuned state of QD devices.
-Thus, even if such specialized algorithms perform well on the development device, they require significant redesigning to be deployable on other QD devices.
-In contrast, when creating BATIS, we focused on developing a platform-agnostic tool that can be applied to a variety of QD devices.
-While the procedural flow of BATIS is predetermined, the executable flow is automatically adjusted based on the general device architecture and connectivity to the measurement instruments, both provided within the configuration files.
-In other words, rather than being programmed to follow a fixed sequence of measurements, BATIS is designed to choose which steps to take and change the expected outcome of each step – depending on the device used.
-Further adjustments are automatically made as the tuning progresses and the algorithm learns the device.
-The device characterization and tuning processes are a repetitive yet essential part of initializing QD-based qubit chips.
-The majority of autotuning efforts have focused on the latter phases of automatic device tuning, including coarse tuning, charge-state tuning, virtualization, and fine-tuning, which is sufficient for quantum dots in a one-dimensional channel or when the initial step is performed by hand [25].
-In modern chips, the scale of the tune-up will soon preclude hand-tuning as fully 2D structures become ubiquitous.
-BATIS addresses the challenges that are about to become limiting in this widespread class of quantum semiconductor chips.
+These channels are not separate until pinch-off voltages, i.e., the gate voltage threshold at which carrier accumulation is eliminated, are determined 111For a given current curve, the pinch-off and the turn-on voltage are mathematically indistinguishable.
+However, when measuring a physical gate, a hysteresis (i.e., a shift of the current curve that depends on the direction of the sweep) can be observed..
+Moreover, since the accumulation and depletion of carriers must occur in alternating regions separated by only tens of nanometers, these channels must interact to enable the formation of QDs separated by tunnel barriers and proximal charge sensors.
+Here, we describe a bootstrapping, autonomous testing, and initialization system (BATIS) that can automatically manipulate gate voltages to ensure that the QD chip is properly accumulated and all required pinch-off voltages are determined.
+BATIS takes as input the device gate layout and a high-level understanding of the desired operation, including (critically) where current channels should be continuous and which gates should divide one channel from another.
+The outputs of BATIS are the gate voltages that make the QD chip functional, with all current channels well delineated and all pinch-off voltages known.
+Importantly, in this work, the sample is kept at around —about 10 times warmer than other typical devices [10, 11]—which is highly desirable for high-temperature operation [12].
+However, higher temperatures broaden key measurement features, such as Coulomb blockade peaks, necessitating careful fitting procedures to extract the desired parameters [13].
+We show that BATIS works well at these high temperatures, which is important both for scalability and automated testing, e.g., using a cryogenic wafer prober [14, 15].
+Previous work on autotuning has shown that correctly executed bootstrapping of QD devices sets the stage for the more specialized tuning phases.
+For example, setting the device topology in a modulation-doped QD device with predefined current channels has been demonstrated in Refs.
+[16, 17, 18].
+Autotuning has similarly been implemented in an undoped SiMOS QD to achieve single-electron occupation [19].
+For nanowires, which by design have a single current channel, automated tuning into fully formed QDs has been demonstrated, provided device-specific details were preconfigured [20].
+Additionally, for undoped, 1D accumulation-mode QD devices, previous work has enabled progress in tuning a device topology [22, 23] and calibrating the charge state [16, 24].
+Methods for automatically detecting unintentional, spurious QDs have also been proposed [25].
+More recently, virtual-gate control of the electrostatic potential landscape confining holes within a 2D QD array has been demonstrated [26].
+However, autonomous bootstrapping for accumulation-mode QD devices has never been attempted.
+Rather, most proposed tuning algorithms assumed that initial steps, such as device characterization and channel formation, have already been accomplished [27].
+Our approach for BATIS differs from the existing bootstrapping algorithms, which often use efficiency as a figure of merit.
+By focusing on speed, such algorithms tend to be specialized to a particular device, allowing little room for benchmarking on different devices or in different experimental setups.
+Even if such algorithms perform well in one setting, they require a significant redesign to work in other settings.
+In contrast, we have specifically designed BATIS to be platform-agnostic, so that it may be applied to a variety of QD devices.
+While the procedural flow of BATIS is well-defined, the executable flow is automatically adjusted to suit the device architecture and external control lines, both of which are specified in the configuration file.
+In other words, rather than following a fixed sequence of measurements, BATIS—in real-time—chooses which steps to take and defines the expected outcomes depending on the device used.
+As the tuning progresses and the algorithm learns the device’s characteristics, BATIS applies its knowledge to automatically customize further measurements.
+Finally, we note that existing autotuning schemes focus mainly on the more advanced stages of a tune-up.
+These include coarse tuning, charge-state tuning, virtualization, and fine-tuning, which are sufficient for QDs formed in a preexisting channel, or when the initial bootstrapping steps are performed by hand [27].
+In modern chips, hand-tuning will soon be precluded as QD processors scale up.
+BATIS is designed to autonomously adjust to the QD layout and perform the initial tune-up stages, including checking for fabrication defects and characterizing local and global current flow, all of which are essential for large-scale qubit implementations.
+It thus addresses the challenges that are about to become limiting in this widespread class of quantum semiconductor chips.
 II BATIS Design
-Figure 1(a) shows a gate-defined QD device that illustrates the critical features of accumulation-mode, undoped devices: multiple channels (three white arrows), regions for forming quantum dots (four in the qubit channel, below, and two single-QD charge sensors, above), and reservoir gates to connect these regions to ohmic contacts for current flow.
+Figure 1(a) shows a gate-defined QD device that illustrates the critical features of accumulation-mode, undoped devices: multiple channels (three white arrows), regions for forming QDs (four in the qubit channel, below, and two single-QD charge sensors, above), and reservoir gates to connect these regions to ohmic contacts for current flow.
 This device exemplifies a commonly used fabrication strategy known as overlapping gates [28, 29].
-The screening gates, shown in yellow, are the closest layer of gates to the top gate oxide in the three-layer gate design and are used to isolate each of the channels from each other and from other regions in the chip.
-The second layer (shown in blue and red) includes the reservoir and plunger gates.
-The reservoirs enable the flow of electrons from the ohmic contacts toward the center of the device where QDs will form in the electrostatic potential profile controlled by each plunger gate.
-The final layer consists of barrier gates (grey), which control the tunnel couplings between QDs—both qubit QDs and charge-sensing QDs—and between the outermost QDs and reservoirs.
-By carefully adjusting the voltages applied to each of the gates, the quantum dot qubit chip with charge sensors is made functional [2, 3, 4, 5].
+The screening gates (yellow) are the closest layer of gates to the top gate oxide in the three-layer gate design and screen electric fields to prevent unwanted regions of 2DEG/2DHG, effectively isolating the channels.
+The second layer includes the reservoir (blue) and plunger (red) gates.
+The reservoirs enable the flow of electrons from the ohmic contacts toward the center of the device, where QDs will form in the electrostatic confined potential.
+The chemical potential of each QD is primarily controlled by the corresponding plunger gate.
+The final layer consists of barrier gates (grey), which control the tunnel couplings between QDs—both qubit QDs and charge-sensing QDs—and reservoirs.
+By carefully adjusting the voltages applied to each of the gates, the QD qubit chip with charge sensors is made functional [2, 3, 4, 5].
 BATIS is designed to automate the task of testing and characterizing all gates and forming all necessary 1D current channels for arbitrary-length linear QD arrays.
-The procedural flow of the single button press algorithm is depicted in Fig.
+The procedural flow of the single-button-press algorithm is depicted in Fig.
 1(b).
@@
 Based on that information, it performs full diagnostics and characterization of each gate and determines the proper configuration for the consecutive stage of the tuning processes.
-Each of the six stages in BATIS involves at least one device measurement followed by data analysis to test the functionality of a selected subset of gates, shown in Fig.
+Each of the six stages in BATIS involves at least one device measurement followed by data analysis to extract characteristics of a selected subset of gates, shown in Fig.
 1(b) by gray rounded rectangles.
-The optional illumination stage is executed on an as-needed basis.
+The optional optical illumination stage is executed on an as-needed basis.
 Illumination involves exposing the QD device to 780 light under carefully chosen gate voltages, to modify its state.
 Such exposure shifts the gate voltage response by modifying the amount (and if desired, the sign) of charge trapped at the semiconductor-oxide interface [21].
-The algorithm follows a conditional logic visually represented with arrows of different colors and textures in Fig.
+The algorithm follows conditional logic visually represented with arrows of different colors in Fig.
 1(b).
-The solid gray arrows indicate the procedural step in the sequence.
-Passing a test for a current stage, marked with yellow diamonds in Fig.
-1(b), initiates a successive measurement and assessment stage, as indicated with solid green arrows.
-Failing a test can result in repeated measurements, device illumination, or tuning termination, depending on the stage.
-The measurement schemes, together with idealized and real-world data associated with each stage, are shown in Fig.
+The gray arrows indicate the procedural step in the sequence.
+Passing the test for the current stage, marked with yellow diamonds in Fig.
+1(b), initiates the next stage, as indicated with green arrows.
+Failing a test can result in repeated measurements, device illumination, or termination of tuning, depending on the stage.
+The physical intuition behind each measurement, together with idealized and real-world data associated with each stage, are shown in Fig.
 2.
 Once completed, BATIS produces a device diagnostics file that stores all information about each gate learned during tuning.
-It also pre-configures the QD device for further use.
+It also preconfigures the QD device for further use.
 The details of the flow of the algorithm are discussed in the following sections.
@@
 The two configuration files used as input for BATIS include descriptions of all gates, desirable current channels, and the QD device connections.
-The files include also details about the software configuration necessary to execute BATIS and a list of variables that need to be considered at each stage.
-These variables consist of instrumentation limits to ensure that the algorithm does not search in ranges not supported by the instruments and device limits to prevent the device from breaking during tune-up.
+The files also include details about the software configuration necessary to execute BATIS and a list of variables that need to be considered at each stage.
+These variables consist of instrumentation limits to ensure that the algorithm does not search in ranges not supported by the instruments and device limits to protect the device from having large damaging voltages applied.
 Details included in configuration files allow the linking of abstract concepts about the device components with the language used in BATIS.
-The functionality of gates is determined automatically from the device architecture that the end user inputs, building a gate-lookup table.
+For example, the functionality of gates is determined automatically from the device architecture that the end user inputs, building a gate-lookup table.
 The files also contain a wire map and instrument information detailing the measurement system setup such that BATIS can automatically read and execute all specialized measurements, connecting the gate-lookup table with the measurement instruments.
 This allows BATIS to remain device- and instrument-agnostic.
-BATIS starts by checking the device for leakage to the fridge ground through the oxide.
+BATIS starts by checking the device for leakage to the fridge ground through the oxide or any possible means in the instrumentation setup.
 The first leakage measurement happens before the electron accumulation during the leakage test stage, see Fig.
@@
 2(a).
-The desired result is a high resistance between all the connections, an entirely yellow matrix.
+The desired result is a high resistance between all the connections.
 Defects during the fabrication process can result in leakage.
@@
 1(b), is to determine the characteristic voltage where enough of the 2DEG is accumulated that a significant amount of current begins to flow between the ohmic contacts, effectively turning the device on.
+Determining the turn-on voltages VP relies on fitting to a phenomenological function that mimics the expected behavior of the signal.
+Prior to fitting, the raw data is preprocessed to clean the signal.
+The preprocessing involves cropping the data to avoid non-linearities at the extreme limits of the data acquisition electronics followed by smoothing the data to remove noise from the signal.
+The processed signal is then fit to a function of the form
+where is the center of the signal and describes the slope at . The final turn-on voltage is defined as
+where is a fixed constant specifying how much the gate needs to be adjusted to turn-on in the units of 222The parameter is defined as part of the configuration file and can be adjusted if needed..
+In addition, the saturation of the curve is defined following Eq.
+(2), with .
+In cases where the saturation of the curve is not captured in the measurement, the last point along the sweep axis becomes .
+The first estimation of the maximum allowed voltage applied to each gate is defined as V.
 Idealized successful and failing measurements for this stage are visualized in the left and right panels in the middle row of Fig.
 2(b), respectively.
-The turn-on process in real QD devices is not always straightforward, as even a good device might not turn on properly when first tested.
+The turn-on process in real QD devices is not always straightforward, as even a good QD device might not turn on properly when first tested.
 If the algorithm encounters a problem within global turn-on, it has a recovery pathway through optional illumination, see Fig.
@@
 II.3 Accumulated Leakage Test
-Once the 2DEG is accumulated, the second, accumulated leakage test is initiated.
+Once the 2DEG is accumulated, the accumulated leakage test is initiated.
 This step proceeds analogously to the first leakage check, except that it does not include the ohmic connections.
@@
 II.4 Screening and Reservoir Characterization
-The turn-on characteristic voltage is taken globally, meaning it does not provide gate-specific information.
+The global accumulation voltage does not provide gate-specific information.
 To learn that information, BATIS performs a series of individual gate characterizations.
@@
 1(b).
-With the 2DEG accumulated and the device turned on, BATIS depletes the voltages on individual gates until the current on the corresponding channel is pinched off, as displayed in the right panel in the second row in Fig.
+With the 2DEG globally accumulated using the voltages selected from Sec.
+II.2, BATIS depletes the voltages on individual gates until the current on the corresponding channel is pinched off, as displayed in the left panel in the second row in Fig.
 2(c).
-BATIS uses these pinch-off curves to determine where to place reservoir gates for proper 2DEG accumulation and establish bounds to operate the screening gates.
-The typical failure modes for this stage include too much depletion and too much accumulation.
+BATIS uses these pinch-off curves to determine the operational configuration for reservoir gates that ensures proper 2DEG accumulation and establishes bounds for operating the screening gates.
+The signals for screening and reservoir gates are preprocessed as described in Sec.
+II.2.
+Once cleaned, the derivative of the current vs.
+voltage signal, , is fit to the derivative of the phenomenological function given in Eq.
+(1).
+The pinch-off, guaranteed isolation, and operating voltages are defined following Eq.
+(2), with , , and , respectively.
+The operating voltage for the central screening gate is handled separately as it has to satisfy two constraints: ensuring that all current channels can be formed without merging.
+By default, .
+If at the subsequent state, a failure to form 1D current channels is observed twice, BATIS adjusts this parameter in increments defined as , where is chosen at random.
+The typical failure edge cases for this stage include too much depletion and too much accumulation; see the second row, the right column of Fig.
+2(c).
 If the channel below a gate is extraordinarily depleted, no voltage can be applied to the gate to accumulate it.
-However, since the screening and reservoir characterization stage is only initialized if all current channels turn on properly at the global accumulation stage, the inability to accumulate implies either an issue with the measurement instrument or hysteresis with the gate’s action upon the 2DEG.
-Similarly, pinching off of all current channels is also confirmed at the global accumulation stage, and thus, it is implausible to observe too much accumulation at this stage.
-To test whether the issue is caused by the instrument, the algorithm executes a repeated measurement.
+On the other hand, if the channel is too accumulated, it is not possible to pinch it off.
+However, both types of issues would have been detected at the global accumulation stage, prompting the initiation of device illumination.
+Thus, the no change in current at this stage has to be caused either by an issue with the instrument or a broken gate.
+To check whether this is the case, the algorithm executes a repeated measurement.
 If the error is observed a second time, it indicates an issue with the gate and the algorithm terminates.
 II.5 Forming 1D Current Channels
-Selectively removing 2DEG from certain regions of the device is done to form the 1D current channels.
-It requires depleting the screening gates and accumulating the plungers and barriers (often collectively called finger gates) such that the current flows between the screening gates along a desirable path, see the left panel in Fig.
-2(c).
+The 1D current channels are formed by selectively removing 2DEG from certain regions of the device.
+It requires depleting the screening gates and accumulating the plungers and barriers (often collectively called finger gates) such that the current flows between the screening gates along the desirable path, as seen in the top row of Fig.
+2(d).
 Determining voltages where 1D channels are properly formed is typically done by sweeping two screening gates simultaneously in a 2D sweep while holding the finger gates constant.
-The formation of the channel manifests as a triangular-sloped region (the so-called triangle plots).
+The formation of the channel manifests as a triangular-sloped region (the so-called triangle plots) [31].
 If the finger gates are not accumulated enough, this region will not form.
@@
 Instead of sweeping two screening gates, BATIS takes a 2D voltage scan while sweeping one screening gate against all of the finger gates between that gate and the other screening gate used to form the 1D channel (typically the central screening gate).
-Since BATIS already learned the necessary configuration for the channel isolation with the central screening gate, this information is now utilized to configure the remaining screening gates.
+BATIS uses the channel isolation (VI) and the operating points (VO) to prepare the device for characterizing the remaining screening gates by sweeping from VI past VP.
 The 1D channel formation technique implemented in BATIS requires a single measurement for each pair of gates, which makes it much more efficient than the standard approaches.
-The plot in the right panel in the second row Fig.
+The plot in the left panel in the second row Fig.
 2(d) exhibits three characteristic regions: a no-current region in the bottom left, a screening-gate-dominated region in the bottom right, and a finger-gate-dominated region in the top left.
 BATIS automatically selects an operating point in the top left part of the plot, near the boundary between the current and the no-current regions.
-The automatic analysis of the regions and their respective boundaries is enabled by fitting a synthetic model to the experimental data, similar to the presentation in [29].
-Additional criteria based on the local properties, e.g., estimated resistance and derivatives of the measurement, are used to optimize the selection of the final point among numerous candidates in the operating region.
-If the finger-gate-dominated region does not form, BATIS illuminates the device with the finger gates at a voltage 400 more negative than the most recent illumination.
+The automatic analysis of the regions and their respective boundaries is enabled by fitting a synthetic model to the experimental data, similar to Ref.
+[32].
+Additional criteria based on the local properties, e.g., estimated resistance and derivatives of the measurement, are used to optimize the selection of the final point V for , among numerous candidates in the operating region.
+If the finger-gate-dominated region does not form, as shown in the right panel in the second row of Fig.
+2(d), BATIS illuminates the device with the finger gates voltages set to be more negative than the most recent illumination.
+This adjustment can be found in the configuration file, we used 400 .
 After such an illumination, the entire device response can change, because of cross-talk and the interdependence of the current paths.
@@
 II.6 Finger Gate Characterization
-Once isolated current channels are formed, finger gate pinch-offs can be measured.
-This is done by assuming that only the nearest neighbors can affect the accumulation under a gate via crosstalk when pinching off the channels.
-Ideally, as the voltage on a given finger gate is lowered, the pinch-off curves should exhibit a smooth, sigmoidal trend, eventually cutting off the current in the respective channels.
+Once isolated current channels are formed, finger gate pinch-offs (VP) can be measured.
+Ideally, as the voltage on a given finger gate is lowered, the pinch-off curves should exhibit a smooth, sigmoidal trend consistent with Eq.
+1, eventually cutting off the current in the respective channels following Eq.
+(2) with .
 An idealized representation of such a curve is visualized in the left panel in the second row in Fig.
@@
 We now demonstrate BATIS on a 1D quadruple Si/SixGe1-x QD device, nominally identical to the one shown in Fig.
-1(a).
-The device’s schematic consisting of 24 gates is shown in Fig.
-3(a).
+1(a) with 24 gates.
 The first layer of gates is made up of four screening gates, Si for , shown in Fig.
-3(a) in yellow.
+1(a) in yellow.
 The next layer consists of five reservoirs, Ri for , and six plunger gates, Pi for .
 The reservoirs are shown in Fig.
-3(a) in blue, while the plungers are shown in red.
+1(a) in blue, while the plungers are shown in red.
 The final layer consists of nine barrier gates, Bi for , shown in Fig.
-3(a) in gray.
-The goal of BATIS is to form three current channels in the quad-QD device, as indicated with white arrows in Fig.
+1(a) in gray.
+The goal of BATIS is to form three current channels in the quad-QD device as it learned from the gate layout in the configuration, as indicated with white arrows in Fig.
 3(a).
@@
 1(d).
-In what follows, we explain how BATIS performs the quad-QD device bootstrapping, testing, and characterizations.
+In what follows, we explain how BATIS navigates the execution of the quad-QD device bootstrapping, testing, and characterizations.
 First Leakage Test.
@@
 In the worst case, a total of resistance measurements are necessary to complete the leakage matrix.
-The leakage test needs to be executed only once per cool-down if under fully automatic control.
+The leakage test needs to be executed only once per cooldown if under fully automatic control.
 This is because BATIS has built-in safety checks during voltage choice selection to avoid exceeding safe voltage breakdowns between neighboring gates for all measurements.
+| Voltages | Run 1 | Run 2 | Run 3 | Run 4 | |
+| I1 | VP () | ||||
+| Vmax () | |||||
+| I2 | VP () | ||||
+| Vmax () | |||||
+| I3 | VP () | ||||
+| Vmax () |
 Global Accumulation.
-Once no leakage is confirmed, BATIS performs the global turn-on stage.
+Next, BATIS performs the global accumulation stage.
 In our demonstration, the algorithm executed three rounds of illumination before the observed global turn-on values were determined to be satisfactory, occurring between and .
-During the first two illuminations, all gates were grounded to reset the device after cooldown and verify that no instruments were erratic.
-After that, for the third illumination, BATIS chose to apply voltages to all screening and reservoir gates.
+By design, all gates are grounded during the first two illuminations.
+The first illumination is intended to completely reset the device while the second is to verify that all instruments work as intended.
+Because BATIS still did not find the gates in the desired range, it chose to apply voltages to all screening and reservoir gates for the third illumination.
 The voltages required to achieve turn-ons at around , calculated following the protocol in Ref.
@@
 1(c), where the pinch-offs on all the screening gates, excluding S2, and all of the reservoir gates are colored based on the learned gate voltage.
-| Channel | Run 1 | Run 2 | Run 3 | Run 4 | |
-| characteristics | |||||
-| I1 | VP () | ||||
-| Vmax () | |||||
-| I2 | VP () | ||||
-| Vmax () | |||||
-| I3 | VP () | ||||
-| Vmax () | |||||
 Accumulated Leakage Test.
 With the global accumulation confirmed, BATIS proceeds to the accumulated leakage test.
-A second leakage matrix is determined, following the same strategy as during the first leakage test.
+A second leakage matrix is measured, following the same procedure as during the first leakage test.
 Like previously, BATIS finds that all resistances exceed 25 .
@@
 The voltage values are quite consistent between the sets of reservoir and screening gates intended to form the desired current channels, a feature from the illumination from earlier [21].
-We note that for gate S1, BATIS acquired signal over , i.e., twice as long as for all remaining gates.
+We note that for gate S1, BATIS acquired the signal over a range, i.e., twice as long as compared to all other screening gates.
 This is likely due to poor fit scoring using the reduced voltage scan range, which caused BATIS to perform additional measurements with an increased sweep range to be more confident in the derived pinch-off value.
@@
 3(b) is truncated to be consistent with the remaining measurements for visualization purposes.
-BATIS sets the reservoir gates two standard deviations above the mean of the phenomenological fit it performed on the signal’s gradient.
-This is to ensure that reservoirs will consistently provide electron flow through the channel as we do not want the cross-talk effects to cause drastic changes in the electron accumulation.
-At the same time, if the reservoirs are too accumulated, they can form spurious channels behind the regions of interest.
-To resolve this, the screening gates are set below their pinch-offs in Fig.
-3(b) to prevent current flow through spurious channels 111The derivative of the screening gate pinch-off was fit to a function like .
-The non-central screening gates were set to be before the mean of the derivative..
-The summary of the pinch-off voltages and the operating point selected by BATIS is presented in Table 2.
-The resulting configuration, where the device is in the correct operating point to perform channel accumulation, is visualized in the central plot in Fig.
+The summary of the pinch-off voltages and the operating points selected by BATIS is presented in Table 2.
+A device in this configuration prepared for the channel accumulation stage is shown in Fig.
 1(c).
@@
 The three points are consolidated into one where the voltage color for each gate is the VO indexed by the channel the gate belongs to.
-| Gate | V () | V () | V () |
-| S1 | |||
-| S3 | |||
-| S4 | |||
 As we discuss in Sec.II.5, the algorithm for 1D current channel formation implemented in BATIS does not follow the typical, triangle-plot-based strategy.
@@
 In other words, the triangle plots and the plots utilized in BATIS contain similar information on different plane cuts through the same space.
+| Gate(s) | V () | V () | V () |
+| S1 | |||
+| S3 | |||
+| S4 | |||
 The accumulation plots in Fig.
@@
 This is due to separation with the S2 gate which divides the 2DEG between the upper readout channels and the lower dot channel.
-Finger Gate Characterization.
-The last stage of BATIS involves finger gate characterization.
-This characterization is done individually for all of the finger gates on the QD device.
-The voltage range for each measurement is always the same and predefined in the configuration file.
-However, due to the Coulomb blockade oscillations, the automated search for the starting voltage for some gates is more challenging than for others.
-This can be seen in the raw data pictured in Fig.
-3(d).
-To ensure that the right point is consistently found, BATIS combines thresholding and slope extraction for each plot to ensure that the pinch-off it finds is robust to noise and spurious fluctuations that do not exceed the threshold.
-The resulting pinch-off values of all of these varying curves for all finger gates are presented in Table 4.
-This configuration is used in the final colorized SEM for the ending state of the algorithm shown in Fig.
-1(d).
 | Gate | VP () | Gate | VP () | Gate | VP () |
@@
 | B5 | |||||
+Finger Gate Characterization.
+The last stage of BATIS involves finger gate characterization.
+This characterization is done individually for all of the finger gates on the QD device.
+The voltage range for each measurement is always the same and predefined in the configuration file.
+However, due to the Coulomb blockade oscillations, the automated search for the starting voltage for some gates is more challenging than for others.
+This can be seen in the raw data pictured in Fig.
+3(d).
+The resulting pinch-off values of all of these varying curves for all finger gates are presented in Table 4.
+This configuration is used in the final colorized SEM for the ending state of the algorithm shown in Fig.
+1(d).
 IV Summary and outlook
-In this paper, we proposed and demonstrated BATIS: a fully automatic procedure for bootstrapping, automated testing, and characterization of accumulation mode, undoped QD devices.
-The prior knowledge about the QD device required to initialize BATIS is limited to the heterostructure, the device geometry, and the device limits.
+In this paper, we propose and demonstrate BATIS, a fully autonomous procedure for bootstrapping, automated testing, and characterization of accumulation mode undoped QD devices.
+The prior knowledge about the QD device required to initialize BATIS is limited to the heterostructure, the device geometry, the ohmic biases, and the device limits.
 BATIS navigates through complex high-dimensional gate space to find the final gate configuration in which all desirable current channels are formed following simple physical principles.
-The goal of BATIS is to test the device’s functionality to confirm that all gates can properly pinch off and that the current channels can be formed.
-Testing and characterizing an unknown QD device is essential for identifying suitable global device states that enable the observation of Coulomb blockade.
-Importantly, the single button press testing and initialization is demonstrated at , eliminating the need for a full cooldown just to learn that a device is not functional.
+The goal of BATIS is to test the device’s functionality to confirm that all gates can properly pinch-off and that all current channels can be formed.
+Testing and characterizing an unknown QD device is essential for identifying global device states that are suitable for observing the Coulomb blockade.
+Importantly, the single-button-press testing and initialization process is demonstrated at , eliminating the need for a full cooldown just to learn that a device is not functional.
 This result is particularly important as we envision a hot large-scale spin-based quantum processor.
 These systems will require qubit initialization, relying on a complete and accurate device characterization.
-One could imagine using BATIS to probe the attributes of many devices spread across a wafer or provide device diagnostics on a partition of a large qubit processor without going to temperatures.
-The flexible structure allows BATIS to be scalable to characterize and initialize any accumulation mode linear array of QD and easily adaptable to other QD platforms.
-In particular, BATIS could straightforwardly mimic the behavior of algorithms proposed for nanowires and GaAs/AlxGa1-xAs by simply invoking skip logic on the steps not necessarily supported for these devices.
-However, it is quite unlikely that any previously proposed algorithms could be easily adapted to the Si/SixGe1-x quad-QD device considered in this work.
-Previous studies have demonstrated a device-agnostic algorithm for certain aspects of QD device tuning, such as setting topology [32, 17] and charge state classification [33, 22].
-However, all these algorithms presuppose that device characterization and foundational tuning steps—determining gate functionality, defining the QD channel(s), or locating regions of interest for the coarse tuning regime—have already been performed.
-BATIS fills this gap, making device characterization easily within reach for linear QD array systems and paving the way toward fully automatic control.
-When linked together, all proposed subroutines can form an autonomous system for controlling arbitrary QD devices up to any desired level of characterization, which will accommodate the growth of QD qubit systems on the path to a scalable quantum computer on this platform.
+One could imagine using BATIS to probe the attributes of many devices spread across a wafer or provide device diagnostics on a portion of a large qubit processor without going to temperatures.
+The flexible structure of BATIS allows for scalable characterization and initialization of any accumulation-mode linear array of QDs.
+It is also easily adaptable to other QD platforms.
+In particular, BATIS could straightforwardly mimic the behavior of algorithms proposed for nanowires and GaAs/AlxGa1-xAs by simply invoking skip logic on the steps not necessary for these devices.
+On the other hand, it is quite unlikely that any previously proposed algorithms could be easily adapted to the Si/SixGe1-x quad-QD device considered in this work.
+Previous studies have demonstrated device-agnostic algorithms for certain subsets of QD device-tuning tasks, such as setting QD topologies [33, 17] or charge-state classifications [34, 24].
+However, these algorithms presuppose that device characterization and foundational tuning steps—such as determining gate functionality, defining the QD channel(s), or locating regions of interest for the coarse-tuning regime—have already been performed.
+BATIS fills this gap, bringing device characterization easily within reach for linear QD arrays and paving the way toward fully automatic control.
+When linked together, the proposed subroutines form an autonomous system for controlling arbitrary QD devices up to any desired level of characterization, putting QD qubit systems on the path to scalable quantum computing.
 Acknowledgements.
@@
 Edge for providing one of the Si/SixGe1-x heterostructures used in this work.
-This work supported in part by ARO grant nos.
-W911NF-23-1-0110 and W911NF-17-1-0274.
+This work was supported in part by ARO grant nos.
+W911NF-17-1-0274, W911NF-23-1-0110, and W911NF-24-2-0043.
+The authors gratefully acknowledge the use of facilities and instrumentation in the UW-Madison Wisconsin Center for Nanoscale Technology, which is partially supported by the Wisconsin Materials Research Science and Engineering Center (NSF DMR-2309000) and the University of Wisconsin-Madison.
 The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S.
@@
 R.
-Hulbert, E.
-Acuna, R.
-S.
-Noah, J.
-W.
-Matten, M.
-P.
-Jura, J.
-A.
-Wright, M.
-T.
-Rakher, and M.
-G.
-Borselli, A flexible design platform for si/sige exchange-only qubits with low disorder, Nano Lett.
+Hulbert, et.
+al, A flexible design platform for si/sige exchange-only qubits with low disorder, Nano Lett.
 22, 1443 (2022), pMID: 34806894.
@@
 L.
-Lawrie, L.
-E.
-A.
-Stehouwer, S.
-D.
-Oosterhout, A.
-Sammak, M.
-Friesen, G.
-Scappucci, S.
-L.
-de Snoo, M.
-Rimbach-Russ, F.
-Borsoi, and M.
-Veldhorst, Operating semiconductor quantum processors with hopping spins, Science 385, 447 (2024).
+Lawrie, et.
+al, Operating semiconductor quantum processors with hopping spins, Science 385, 447 (2024).
 - George et al.
@@
 K.
-Harper, D.
-Keith, L.
-Lampert, F.
-Luthi, F.
-A.
-Mohiyaddin, S.
-Murcia, R.
-Nair, R.
-Nahm, A.
-Nethwewala, S.
-Neyens, R.
-D.
-Raharjo, C.
-Rogan, R.
-Savytskyy, T.
-F.
-Watson, J.
-Ziegler, O.
-K.
-Zietz, R.
-Pillarisetty, N.
-C.
-Bishop, S.
-A.
-Bojarski, J.
-Roberts, and J.
-S.
-Clarke, 12-spin-qubit arrays fabricated on a 300 mm semiconductor manufacturing line (2024), arXiv:2410.16583 [cond-mat.mes-hall] .
+Harper, et.
+al, 12-spin-qubit arrays fabricated on a 300 mm semiconductor manufacturing line (2024), arXiv:2410.16583 [cond-mat.mes-hall] .
+- Note [2] For a given current curve, the pinch-off and the turn-on voltage are mathematically indistinguishable.
+However, when measuring a physical gate, a hysteresis (i.e., a shift of the current curve that depends on the direction of the sweep) can be observed.
 - Borsoi et al.
@@
 Wagner, F.
-Borjans, E.
-J.
-Connors, J.
-Corrigan, M.
-J.
-Curry, D.
-Keith, R.
-Kotlyar, L.
-F.
-Lampert, M.
-T.
-Mądzik, K.
-Millard, F.
-A.
-Mohiyaddin, S.
-Pellerano, R.
-Pillarisetty, M.
-Ramsey, R.
-Savytskyy, S.
-Schaal, G.
-Zheng, J.
-Ziegler, N.
-C.
-Bishop, S.
-Bojarski, J.
-Roberts, and J.
-S.
-Clarke, Probing single electrons across 300-mm spin qubit wafers, Nature 629, 80 (2024).
-- Baart et al.
-[2016] T.
-A.
-Baart, P.
-T.
-Eendebak, C.
-Reichl, W.
-Wegscheider, and L.
-M.
-K.
-Vandersypen, Computer-automated tuning of semiconductor double quantum dots into the single-electron regime, Appl.
-Phys.
-Lett.
-108, 213104 (2016).
+Borjans, et.
+al, Probing single electrons across 300-mm spin qubit wafers, Nature 629, 80 (2024).
 - Darulová et al.
@@
 Zwolak, Autonomous bootstrapping of quantum dot devices, arXiv:2407.20061 (2024).
+- Baart et al.
+[2016] T.
+A.
+Baart, P.
+T.
+Eendebak, C.
+Reichl, W.
+Wegscheider, and L.
+M.
+K.
+Vandersypen, Computer-automated tuning of semiconductor double quantum dots into the single-electron regime, Appl.
+Phys.
+Lett.
+108, 213104 (2016).
 - Lapointe-Major et al.
@@
 Isella, M.
-De Kruijf, M.
-J.
-Carballido, S.
-Svab, A.
-V.
-Kuhlmann, S.
-Geyer, F.
-N.
-M.
-Froning, H.
-Moon, M.
-A.
-Osborne, D.
-Sejdinovic, G.
-Katsaros, D.
-M.
-Zumbühl, G.
-A.
-D.
-Briggs, and N.
-Ares, Cross-architecture tuning of silicon and SiGe-based quantum devices using machine learning, Sci.
+De Kruijf, et.
+al, Cross-architecture tuning of silicon and SiGe-based quantum devices using machine learning, Sci.
 Rep.
 14, 17281 (2024).
+- Wolfe et al.
+[2024] M.
+A.
+Wolfe, B.
+X.
+Coe, J.
+S.
+Edwards, T.
+J.
+Kovach, T.
+McJunkin, B.
+Harpt, D.
+Savage, M.
+G.
+Lagally, R.
+McDermott, M.
+Friesen, et.
+al, Control of threshold voltages in Si/Si0.7 Ge0.3 quantum devices via optical illumination, Physical Review Applied 22, 034044 (2024).
 - Zwolak et al.
@@
 - Zwolak and Taylor [2023] J. P. Zwolak and J. M. Taylor, Colloquium: Advances in automation of quantum dot devices control, Rev. Mod. Phys. 95, 011006 (2023).
-- Wolfe et al.
-[2024] M.
-A.
-Wolfe, B.
-X.
-Coe, J.
-S.
-Edwards, T.
-J.
-Kovach, T.
-McJunkin, B.
-Harpt, D.
-Savage, M.
-G.
-Lagally, R.
-McDermott, M.
-Friesen, S.
-Kolkowitz, and M.
-A.
-Eriksson, Control of threshold voltages in Si/Si0.7 Ge0.3 quantum devices via optical illumination, Physical Review Applied 22, 034044 (2024).
 - Zajac et al.
@@
 Eriksson, Fabrication process and failure analysis for robust quantum dots in silicon, Nanotechnology 31, 505001 (2020).
+- Note [3] The parameter is defined as part of the configuration file and can be adjusted if needed.
+- McJunkin [2021] T.
+W.
+McJunkin, Heterostructure Modifications, Fabrication Improvements, and Measurements Automation of Si/SiGe Quantum Dots for Quantum Computation, Ph.D.
+thesis, The University of Wisconsin-Madison, Madison, WI, USA (2021).
 - Schug et al.
@@
 Zwolak, Automation of Quantum Dot Measurement Analysis via Explainable Machine Learning (2024), version Number: 4.
-- Note [2] The derivative of the screening gate pinch-off was fit to a function like . The non-central screening gates were set to be before the mean of the derivative.
-- McJunkin [2021] T.
-W.
-McJunkin, Heterostructure Modifications, Fabrication Improvements, and Measurements Automation of Si/SiGe Quantum Dots for Quantum Computation, Ph.D.
-thesis, The University of Wisconsin-Madison, Madison, WI, USA (2021).
 - Ziegler et al.

### v3

@@
-BATIS: Bootstrapping, Autonomous Testing, and Initialization System for Quantum Dot Devices
+Bootstrapping and autonomous tuning system for Si/SiGe multi-quantum dot devices
 Abstract
 Semiconductor quantum dot (QD) devices have become central to advancements in spin-based quantum computing.
-As the complexity of QD devices grows, manual tuning becomes increasingly infeasible, necessitating robust and scalable autotuning solutions.
-Tuning large arrays of QD qubits depends on efficient choices of automated protocols.
-Here, we introduce a bootstrapping, autonomous testing, and initialization system (BATIS) designed to streamline QD device evaluation and calibration.
-BATIS navigates high-dimensional gate voltage spaces, automating essential steps such as leakage testing and gate characterization.
-For forming the current channels, BATIS follows a non-standard approach that requires a single measurement regardless of the number of channels.
+However, the increasing complexity of modern QD devices makes calibration and control—particularly at elevated temperatures—a bottleneck to progress, highlighting the need for robust and scalable autonomous solutions.
+A major hurdle arises from trapped charges within the oxide layers, which induce random offset voltage shifts on gate electrodes, with a standard deviation of approximately 83 of variation within state-of-the-art present-day devices.
+Efficient characterization and tuning of large arrays of QD qubits depend on choices of automated protocols.
+Here, we introduce a physically intuitive framework for a bootstrapping, autonomous testing, and initialization system (BATIS) designed to streamline QD device evaluation and calibration.
+BATIS navigates high-dimensional gate voltage spaces, automating essential steps such as leakage testing, formation of all current channels, and gate characterization in the presence of trapped charges.
+For forming the current channels, BATIS follows a non-standard approach that requires a single set of measurements regardless of the number of channels.
 Demonstrated at on a quad-QD Si/SixGe1-x device, BATIS eliminates the need for deep cryogenic environments during initial device diagnostics, significantly enhancing scalability and reducing setup times.
@@
 In recent years, the number of QDs in such devices has risen rapidly [6, 7, 8].
-In the near term, it will no longer be possible to tune up many dots simultaneously using human-based techniques.
-The field of autotuning has arisen to address this challenge.
-Modern semiconductor QD qubit devices have two critical features that enhance their performance, while simultaneously making autotuning challenging and essential.
-First, the active region of the device—the QDs, charge sensors, and reservoirs—contains no doping.
-All charge carriers—electrons or holes—are induced by gate voltages.
-No current can flow until gate voltages are applied and carriers are accumulated, yet turn-on voltages are unknown a priori.
-Furthermore, because of device drift and hysteresis, the exact voltage configuration changes not only between cooldowns but even between successive initialization steps.
+In the near term, it will no longer be possible to tune up QD qubit arrays manually.
+The field of autotuning has arisen to address this challenge [9, 10].
+Modern semiconductor QD qubit devices have two critical features that enhance their performance while simultaneously making autotuning challenging and essential.
+First, the active regions of the device—the qubit QDs, charge sensors, and reservoirs—contain no doping.
+All charge carriers—electrons or holes—are induced by gate voltages and no current can flow until gate voltages are applied and carriers are accumulated.
+At the same time, the charges trapped within the gate-oxide heterostructure of the QD devices make it impossible to know a priori the precise turn-on voltages, i.e., the gate voltage threshold at which carrier accumulation begins.
+The device drift and hysteresis (i.e., a shift of the current curve that depends on the direction of the sweep) introduce additional variability in voltage configuration not only between cooldowns but even between successive initialization steps.
 Second, QD device chips have multiple current channels that often serve different purposes.
-Some can host an array of QD qubits, while others provide a home for charge sensors.
-These channels are not separate until pinch-off voltages, i.e., the gate voltage threshold at which carrier accumulation is eliminated, are determined 111For a given current curve, the pinch-off and the turn-on voltage are mathematically indistinguishable.
-However, when measuring a physical gate, a hysteresis (i.e., a shift of the current curve that depends on the direction of the sweep) can be observed..
-Moreover, since the accumulation and depletion of carriers must occur in alternating regions separated by only tens of nanometers, these channels must interact to enable the formation of QDs separated by tunnel barriers and proximal charge sensors.
-Here, we describe a bootstrapping, autonomous testing, and initialization system (BATIS) that can automatically manipulate gate voltages to ensure that the QD chip is properly accumulated and all required pinch-off voltages are determined.
+Some host an array of QD qubits, while others provide a home for charge sensors.
+These channels are not separated until pinch-off voltages, i.e., the gate voltage threshold at which carrier accumulation is eliminated, are determined 111Mathematically, the pinch-off and the turn-on voltage are indistinguishable for a given current curve.
+However, when measuring a physical gate, a hysteresis can be observed..
+The size of the QD introduces additional challenges: the accumulation and depletion of carriers must occur in alternating regions separated by only tens of nanometers.
+While spatially separated, the channels must remain coupled to ensure proper operation of proximal charge sensors on nearby QDs.
+In this work, we introduce and demonstrate an intuitive bootstrapping, autonomous testing, and initialization system (BATIS) that can automatically manipulate gate voltages to ensure proper formation of all current channels.
 BATIS takes as input the device gate layout and a high-level understanding of the desired operation, including (critically) where current channels should be continuous and which gates should divide one channel from another.
-The outputs of BATIS are the gate voltages that make the QD chip functional, with all current channels well delineated and all pinch-off voltages known.
-Importantly, in this work, the sample is kept at around —about 10 times warmer than other typical devices [10, 11]—which is highly desirable for high-temperature operation [12].
-However, higher temperatures broaden key measurement features, such as Coulomb blockade peaks, necessitating careful fitting procedures to extract the desired parameters [13].
-We show that BATIS works well at these high temperatures, which is important both for scalability and automated testing, e.g., using a cryogenic wafer prober [14, 15].
-Previous work on autotuning has shown that correctly executed bootstrapping of QD devices sets the stage for the more specialized tuning phases.
-For example, setting the device topology in a modulation-doped QD device with predefined current channels has been demonstrated in Refs.
-[16, 17, 18].
-Autotuning has similarly been implemented in an undoped SiMOS QD to achieve single-electron occupation [19].
-For nanowires, which by design have a single current channel, automated tuning into fully formed QDs has been demonstrated, provided device-specific details were preconfigured [20].
-Additionally, for undoped, 1D accumulation-mode QD devices, previous work has enabled progress in tuning a device topology [22, 23] and calibrating the charge state [16, 24].
+The output of BATIS is a comprehensive device characterization, specifying the gate voltage settings needed for QD chip operation, with clearly delineated current channels and identified pinch-off voltages.
+Previous work on autotuning has shown that correctly executed bootstrapping of QD devices sets the stage for the more specialized tuning phases [9].
+For example, setting the device topology has been demonstrated for undoped accumulation-mode QD devices [12, 13].
+Automated calibration of the charge state has been demonstrated for a modulation-doped QD device [14], an undoped SiMOS QD device [15], and undoped accumulation-mode QD devices [16].
 Methods for automatically detecting unintentional, spurious QDs have also been proposed [17].
 More recently, virtual-gate control of the electrostatic potential landscape confining holes within a 2D QD array has been demonstrated [18].
-However, autonomous bootstrapping for accumulation-mode QD devices has never been attempted.
-Rather, most proposed tuning algorithms assumed that initial steps, such as device characterization and channel formation, have already been accomplished [27].
-Our approach for BATIS differs from the existing bootstrapping algorithms, which often use efficiency as a figure of merit.
-By focusing on speed, such algorithms tend to be specialized to a particular device, allowing little room for benchmarking on different devices or in different experimental setups.
-Even if such algorithms perform well in one setting, they require a significant redesign to work in other settings.
-In contrast, we have specifically designed BATIS to be platform-agnostic, so that it may be applied to a variety of QD devices.
-While the procedural flow of BATIS is well-defined, the executable flow is automatically adjusted to suit the device architecture and external control lines, both of which are specified in the configuration file.
-In other words, rather than following a fixed sequence of measurements, BATIS—in real-time—chooses which steps to take and defines the expected outcomes depending on the device used.
-As the tuning progresses and the algorithm learns the device’s characteristics, BATIS applies its knowledge to automatically customize further measurements.
-Finally, we note that existing autotuning schemes focus mainly on the more advanced stages of a tune-up.
-These include coarse tuning, charge-state tuning, virtualization, and fine-tuning, which are sufficient for QDs formed in a preexisting channel, or when the initial bootstrapping steps are performed by hand [27].
-In modern chips, hand-tuning will soon be precluded as QD processors scale up.
-BATIS is designed to autonomously adjust to the QD layout and perform the initial tune-up stages, including checking for fabrication defects and characterizing local and global current flow, all of which are essential for large-scale qubit implementations.
-It thus addresses the challenges that are about to become limiting in this widespread class of quantum semiconductor chips.
-II BATIS Design
-Figure 1(a) shows a gate-defined QD device that illustrates the critical features of accumulation-mode, undoped devices: multiple channels (three white arrows), regions for forming QDs (four in the qubit channel, below, and two single-QD charge sensors, above), and reservoir gates to connect these regions to ohmic contacts for current flow.
+Automated tuning of QD devices from a coldstart has been demonstrated for modulation-doped QD devices with predefined current channels [19, 20, 21] as well as for SiFinFETs and nanowires, which by design have a single current channel [21].
+However, all these demonstrations assumed a prior knowledge of the device and a non-negligible device-specific preconfiguration.
+The first fully autonomous bootstrapping algorithm, intended for initiation on a pristine modulation-doped QD device, was introduced in Ref.
+[22].
+However, the proposed algorithm is platform-specific and assumes the QD device to be functional.
+Autonomous characterization and bootstrapping for accumulation-mode QD devices have never been attempted.
+BATIS addresses a key challenge in bootstrapping the QD devices: the diversity of device architectures and control layouts.
+All initialization tools proposed to date rely on rigid measurement sequences tailored to specific device designs, limiting their applicability across platforms.
+BATIS, in contrast, is designed to be platform-agnostic and thus can operate across a wide range of QD devices by dynamically adapting its procedures.
+While the procedural flow of BATIS is well-defined at a high level, the executable flow—the actual sequence of measurements and operations—is automatically adjusted in real time based on the device’s configuration and control wiring, as specified in a flexible configuration file.
+In other words, throughout execution, BATIS adjusts which steps to take and what outcomes to expect based on the device’s observed behavior.
+As tuning progresses, BATIS further tailors subsequent measurements to match the discovered characteristics of the system, ensuring efficient and device-specific calibration without manual intervention.
+Importantly, the sample used to demonstrate BATIS is kept at around 1.3 , which is about 10 times warmer than other typical devices [23, 24].
+High temperature operation is critical for practical quantum computer implementations [25].
+However, higher temperatures broaden key measurement features, such as Coulomb blockade peaks, necessitating careful fitting procedures to extract the desired parameters [26].
+The industry has demonstrated promising efforts in characterizing the performance of industry-manufactured spin qubit devices at 1.6 using the cryogenic 300- wafer prober [27, 28].
+Autotuning of a single-electron transistor (SET) charge sensor in Si at 1.5 has also been demonstrated [29].
+We demonstrate that BATIS performs reliably at temperatures above 1.0 and in the presence of trapped charges.
+It autonomously adapts to the specific QD layout and carries out critical initial tune-up stages, including identifying fabrication defects and characterizing both local and global current flow.
+These capabilities directly address emerging bottlenecks in scaling quantum semiconductor chips, making BATIS a powerful tool for enabling large-scale qubit implementations.
+This paper is organized as follows: In Sec.
+II, we motivate the necessity of autotuning from the perspective of trapped charges.
+The bootstrap tuning procedure is described in detail in Sec.
+III.
+The experimental demonstration of BATIS to fully autonomously tune-up an accumulation mode Si/SixGe1-x QD device at 1.3 is described in Sec.
+IV.
+We conclude with a discussion of future directions in Sec.
+V.
+II Trapped Charge: A Barrier to Robust Control
+A primary challenge in automating current channel formation in QD devices stems from trapped charges within the QD array, which introduce substantial variability in the required gate voltages.
+Charge trapping refers to the immobilization of carriers at unknown locations due to, e.g., structural imperfections and defects on surfaces and interfaces of the heterostructure hosting the QDs.
+Stray charges can accumulate in any region of the heterostructure, including at buried interfaces, in the oxide layer at the top of the stack, and in particular, at the semiconductor-oxide interface, where there is an extremely high density of interface states [30].
+Trapped charges in QD arrays present multiple challenges that degrade device performance [31].
+First, they reduce the effective carrier concentration, which impedes transport and reduces mobility.
+In addition, trapped charges introduce temporal fluctuations in the local electric field, resulting in increased charge noise that compromises the stability and coherence of quantum states within the QDs.
+Over time, charge trapping can also lead to electrostatic disorder at the interface between the QD array and surrounding materials, further complicating device tunability and limiting the precision required for quantum control.
+While trapped charges are largely locked into place as the device is cooled down from room temperature, they can nevertheless slowly drift over time, even at temperatures [32].
+Moreover, these processes are not well controlled, leading to nonuniform charge accumulation that can vary from one cooldown to the next.
+As a result, significant and nonuniform voltage shifts may be required to offset the effects of trapped charges, and to achieve uniform gate operation across a device, even for nominally identical gate electrodes within a single device.
+Such variability is undesirable for large-scale applications, and several techniques have been developed to mitigate this problem [33].
+One such approach involves a simple “reset” of the trapped charges by optically illuminating the device in situ using a laser diode [34].
+Illumination temporarily excites electrons above the bandgap of the host semiconductor, enabling the redistribution of immobile, unwanted trapped charges that otherwise remain fixed at low temperatures.
+Unfortunately, this procedure is not guaranteed to improve device performance and often needs to be repeated several times through trial and error, with different applied gate voltages.
+The effect of trapped charges can be illustrated by simulating channel formation in a typical QD device.
+Figures 1(a-b) show a gate-defined, accumulation-mode, undoped QD device with multiple current channels, six designated regions for forming QDs (under the six plunger gates, shown in red), and reservoir gates (shown in blue) to connect these regions to ohmic contacts for current flow.
 This device exemplifies a commonly used fabrication strategy known as overlapping gates [35, 36].
-The screening gates (yellow) are the closest layer of gates to the top gate oxide in the three-layer gate design and screen electric fields to prevent unwanted regions of 2DEG/2DHG, effectively isolating the channels.
-The second layer includes the reservoir (blue) and plunger (red) gates.
-The reservoirs enable the flow of electrons from the ohmic contacts toward the center of the device, where QDs will form in the electrostatic confined potential.
-The chemical potential of each QD is primarily controlled by the corresponding plunger gate.
-The final layer consists of barrier gates (grey), which control the tunnel couplings between QDs—both qubit QDs and charge-sensing QDs—and reservoirs.
+The screening gates (shown in yellow) form the bottommost layer in the three-layer gate architecture, positioned closest to the gate oxide.
+Their primary role is to screen electric fields, thereby preventing the unintentional formation of 2DEG or 2DHG regions and electrically isolating the intended transport channels.
+The second layer includes the reservoir and plunger gates.
+The reservoirs enable the flow of electrons from the ohmic contacts toward the center of the device, where QDs will form in the electrostatic confinement potential.
+Each plunger gate primarily controls the chemical potential of the QD below it.
+The final layer consists of barrier gates (shown in gray), which control the tunnel couplings between QDs—both qubit QDs and charge-sensing QDs—and reservoirs.
 By carefully adjusting the voltages applied to each of the gates, the QD qubit chip with charge sensors is made functional [2, 3, 4, 5].
-BATIS is designed to automate the task of testing and characterizing all gates and forming all necessary 1D current channels for arbitrary-length linear QD arrays.
-The procedural flow of the single-button-press algorithm is depicted in Fig.
+To model the effects of trapped charge, we introduce randomly varying charge densities on a tiled grid array at the oxide-semiconductor interface, with the density of each tile selected at random from a normal distribution with the standard deviation corresponding to a shift magnitude of , consistent with experimental observations for silicon QDs [28].
+Each tile is , ensuring that all gates encompass multiple regions of randomly selected charge densities.
+In the case where no random trapped charge is included, the ideal charge densities in the quantum well form coherent channels halfway between the screening gates, as indicated with the arrows in Fig.
 1(b).
-BATIS takes as input two configuration files describing the device design and certain generic characteristics of the device and instruments.
-Based on that information, it performs full diagnostics and characterization of each gate and determines the proper configuration for the consecutive stage of the tuning processes.
-Each of the six stages in BATIS involves at least one device measurement followed by data analysis to extract characteristics of a selected subset of gates, shown in Fig.
-1(b) by gray rounded rectangles.
-The optional optical illumination stage is executed on an as-needed basis.
-Illumination involves exposing the QD device to 780 light under carefully chosen gate voltages, to modify its state.
-Such exposure shifts the gate voltage response by modifying the amount (and if desired, the sign) of charge trapped at the semiconductor-oxide interface [21].
-The algorithm follows conditional logic visually represented with arrows of different colors in Fig.
-1(b).
-The gray arrows indicate the procedural step in the sequence.
-Passing the test for the current stage, marked with yellow diamonds in Fig.
-1(b), initiates the next stage, as indicated with green arrows.
-Failing a test can result in repeated measurements, device illumination, or termination of tuning, depending on the stage.
-The physical intuition behind each measurement, together with idealized and real-world data associated with each stage, are shown in Fig.
+In contrast, Figs.
+1(c) and 1(e) show results obtained when randomly distributed trapped charges are included, resulting in current channels that flow beneath screening gates and pinched channels, respectively.
+Both problems can be resolved by careful adjustment of the various gate voltages to recover the coherent channels shown in Figs.
+1(d) and 1(f), respectively.
+While misdirected and pinched channels are easy to identify in simulations, they are much more difficult to diagnose in experiments.
+The issue of current flowing beneath the screening gate and unintentionally connecting multiple channels, as illustrated in Fig.
+1(c), is intrinsic to 2D gate-defined systems and does not exist in 1D nanowire and nanotube devices [38, 39].
+As such, it is of particular interest in this work.
+The recovery of misdirected current to form the desired coherent channels requires the simultaneous tuning of multiple gates.
+The protocols developed in this work are specially designed to address and characterize such 2D challenges.
+III BATIS: Design and implementation
+BATIS is designed to automate the testing, characterization, and initialization of all gates and the forming of all necessary 1D current channels for linear QD arrays.
+The procedural flow of the single-button-press algorithm, depicted in Fig.
+2, is rooted in the underlying device physics, making it both interpretable and robust, while also enabling adaptability across devices and scalability to arbitrary-length 1D arrays.
+BATIS takes as input two configuration files describing the device design and certain generic specifications of the device and instruments.
+These files include detailed descriptions of all gates, the desired current channels, and the QD device connectivity.
+Additionally, they define a wire map of the experimental setup and a set of variables associated with each stage of the tuning process.
+These variables include instrumentation constraints, which prevent the algorithm from exploring unsupported operating ranges, and device safety limits, which protect the hardware from excessive or damaging voltages.
+Using this information, BATIS performs comprehensive diagnostics and characterization of each gate and automatically determines the appropriate configuration for the next stage of the tuning sequence, as depicted in Fig.
 2.
-Once completed, BATIS produces a device diagnostics file that stores all information about each gate learned during tuning.
+The configuration files establish a link between abstract descriptions of device components and the operational language used within BATIS.
+For instance, the functionality of each gate is automatically inferred from the device architecture provided by the user, enabling BATIS to construct a gate lookup table.
+When combined with the wire map, this allows BATIS to automatically identify and perform the appropriate specialized measurements by linking the logical gate definitions to the physical measurement instruments.
+This abstraction layer ensures that BATIS remains agnostic to both the specific device layout and the instrumentation used.
+The algorithm follows conditional logic, visually represented with boxes and arrows of different colors in Fig.
+2.
+Each of the six stages in BATIS, indicated in Fig.
+2 with gray rounded rectangles, involves at least one optimization of a device measurement or manipulation followed by data analysis to extract characteristics of a selected subset of gates.
+The measurements typical of each stage are detailed in Fig.
+3, with the top row in each panel explaining the measurement, and the middle and bottom rows showing for each step a cartoon representation of the data (in each case, showing success in the left and failure on the right) and an example of measured experimental data, respectively.
+The procedural steps in BATIS are indicated with gray arrows.
+Passing the functionality test for the current stage, marked with yellow diamonds in Fig.
+2, initiates the next stage, as indicated with green arrows.
+Failing a test, indicated with red arrows, results in repeated measurements, device optical illumination to adjust the trapped charges, or bootstrapping termination, depending on the issues revealed by BATIS and the conclusions it draws based on the device physics.
+Once completed, BATIS produces a device diagnostic file that stores all information about each gate learned during tuning.
+When executed successfully, it verifies that all gates are functioning as intended and that isolated regions can be formed in well-defined 1D channels beneath the designated plunger gates.
 It also preconfigures the QD device for further use.
-The details of the flow of the algorithm are discussed in the following sections.
-II.1 Initialization and First Leakage Test
-The two configuration files used as input for BATIS include descriptions of all gates, desirable current channels, and the QD device connections.
-The files also include details about the software configuration necessary to execute BATIS and a list of variables that need to be considered at each stage.
-These variables consist of instrumentation limits to ensure that the algorithm does not search in ranges not supported by the instruments and device limits to protect the device from having large damaging voltages applied.
-Details included in configuration files allow the linking of abstract concepts about the device components with the language used in BATIS.
-For example, the functionality of gates is determined automatically from the device architecture that the end user inputs, building a gate-lookup table.
-The files also contain a wire map and instrument information detailing the measurement system setup such that BATIS can automatically read and execute all specialized measurements, connecting the gate-lookup table with the measurement instruments.
-This allows BATIS to remain device- and instrument-agnostic.
-BATIS starts by checking the device for leakage to the fridge ground through the oxide or any possible means in the instrumentation setup.
-The first leakage measurement happens before the electron accumulation during the leakage test stage, see Fig.
-1(b).
-The resistances of each connection are measured with respect to each other and compiled into a leakage matrix, as shown in the second and third row of Fig.
-2(a).
+In the following, we explain each state of the algorithm.
+III.1 First Leakage Test
+When initiated, BATIS first checks the device for leakage to the fridge ground through the oxide or any other possible means in the instrumentation setup.
+Leakage can be caused by, e.g., defects arising during the fabrication or packaging processes.
+The first leakage measurement happens before the electron accumulation, during the leakage test stage; see Fig.
+2.
+The resistances of all the connections are measured with respect to each other and fridge ground, and they are compiled into a leakage matrix, as exemplified in the second and third row in Fig.
+3(a).
+Leakage manifests as off-colored cells in the matrix, as shown in the right panel in the second row and in the third row of Fig.
+3(a).
 The desired result is a high resistance between all the connections.
-Defects during the fabrication process can result in leakage.
-Leakage manifests as off-colored cells in the matrix, as shown in the right panel in the second row and in the third row of Fig.
-2(a).
 Should leakage be detected, BATIS terminates and flags the device as unusable.
-II.2 Global Accumulation
-For accumulation mode QD devices no 2DEG is present at zero applied gate voltage.
-Applying a non-zero bias lowers the conduction band minimum below the Fermi level, accumulating a 2DEG in the quantum well near the Si/SixGe1-x interface.
-The goal of the global accumulation stage, see Fig.
-1(b), is to determine the characteristic voltage where enough of the 2DEG is accumulated that a significant amount of current begins to flow between the ohmic contacts, effectively turning the device on.
-Determining the turn-on voltages VP relies on fitting to a phenomenological function that mimics the expected behavior of the signal.
-Prior to fitting, the raw data is preprocessed to clean the signal.
-The preprocessing involves cropping the data to avoid non-linearities at the extreme limits of the data acquisition electronics followed by smoothing the data to remove noise from the signal.
-The processed signal is then fit to a function of the form
-where is the center of the signal and describes the slope at . The final turn-on voltage is defined as
-where is a fixed constant specifying how much the gate needs to be adjusted to turn-on in the units of 222The parameter is defined as part of the configuration file and can be adjusted if needed..
-In addition, the saturation of the curve is defined following Eq.
-(2), with .
-In cases where the saturation of the curve is not captured in the measurement, the last point along the sweep axis becomes .
-The first estimation of the maximum allowed voltage applied to each gate is defined as V.
+III.2 Global Accumulation
+For accumulation mode QD devices, no 2DEG is present at zero applied gate voltage.
+Applying a non-zero positive bias lowers the conduction band minimum below the Fermi level, accumulating a 2DEG in the quantum well near the Si/SixGe1-x interface.
+The goal of the global accumulation stage is to determine the characteristic turn-on voltage where enough of the 2DEG is accumulated that a significant amount of current begins to flow between the ohmic contacts, effectively turning the device on; see Fig.
+3(b).
+Since BATIS assumes no a priori knowledge about the device response, is determined through an iterative approach, with increased incrementally until BATIS is confident that is observed in the measured window.
 Idealized successful and failing measurements for this stage are visualized in the left and right panels in the middle row of Fig.
-2(b), respectively.
-The turn-on process in real QD devices is not always straightforward, as even a good QD device might not turn on properly when first tested.
-If the algorithm encounters a problem within global turn-on, it has a recovery pathway through optional illumination, see Fig.
-1(b).
-Illumination is initiated when the device fails to turn on or the turn-on voltage is much too high to be usable.
+3(b), respectively.
+The turn-on process in real QD devices is not always straightforward, as even fully functional devices may fail to turn on correctly during initial testing.
+If the algorithm encounters a problem during global turn-on, it has a recovery pathway through optional illumination via a 780 light to adjust the microscopic configuration, magnitude, and sign of trapped charge, which can dramatically alter the device’s response to gate voltages [34]; see Fig.
+2.
 After illumination, the global accumulation stage is re-initiated.
@@
 Whether or not illumination should be used is an option that is pre-defined in the configuration file.
-II.3 Accumulated Leakage Test
-Once the 2DEG is accumulated, the accumulated leakage test is initiated.
+Determining the turn-on voltages VP relies on fitting to a phenomenological function that mimics the expected behavior of the signal.
+Prior to fitting, the raw data is preprocessed to clean the signal.
+The preprocessing involves cropping the data to avoid non-linearities at the extreme limits of the data acquisition electronics, followed by smoothing the data to remove noise from the signal and normalization.
+We then parameterize the response of the device using
+where is the center of the step function and is the slope at .
+This function mimics the expected exponential behavior of the Fermi-Dirac distribution of electron energies with respect to gate voltages [40, 41, 42].
+Similarly to the characteristics of the metal–oxide–semiconductor field-effect transistors (MOSFETs), an expected saturation will occur, leading to a roughly sigmoidal shape [43].
+The final turn-on voltage is recorded as
+where is a fixed constant specifying how much the gate needs to be adjusted to turn-on in the units of 222The parameter is defined as part of the configuration file and can be adjusted if needed..
+The safe maximum allowed voltage applied to each gate, , is then updated to better match the saturation of the curve, such that it is less than or equal to the original.
+This voltage is later used as the starting point for the characterization of the screening and reservoir pinch-offs.
+III.3 Accumulated Leakage Test
+With the global accumulation confirmed, BATIS proceeds to the accumulated leakage test.
 This step proceeds analogously to the first leakage check, except that it does not include the ohmic connections.
@@
 Should that occur, BATIS terminates.
-II.4 Screening and Reservoir Characterization
+III.4 Screening and Reservoir Characterization
 The global accumulation voltage does not provide gate-specific information.
 To learn that information, BATIS performs a series of individual gate characterizations.
-The first of these is performed at the screening and reservoir characterization stage in Fig.
-1(b).
-With the 2DEG globally accumulated using the voltages selected from Sec.
-II.2, BATIS depletes the voltages on individual gates until the current on the corresponding channel is pinched off, as displayed in the left panel in the second row in Fig.
-2(c).
-BATIS uses these pinch-off curves to determine the operational configuration for reservoir gates that ensures proper 2DEG accumulation and establishes bounds for operating the screening gates.
-The signals for screening and reservoir gates are preprocessed as described in Sec.
-II.2.
-Once cleaned, the derivative of the current vs.
-voltage signal, , is fit to the derivative of the phenomenological function given in Eq.
-(1).
+The first of these is performed at the screening and reservoir characterization stage; see Fig.
+2.
+With the 2DEG globally accumulated, BATIS depletes the voltages on individual gates until the current on the corresponding channel is pinched off, as displayed in the left panel in the second row in Fig.
+3(c).
+BATIS uses these pinch-off curves to determine the operating points for reservoir gates, ensuring proper 2DEG accumulation, and to determine the operating bounds for the screening gates.
+The derivative of the response signals for screening and reservoir gate pinch-offs are preprocessed in a similar fashion as described in Sec.
+III.2.
 The pinch-off, guaranteed isolation, and operating voltages are defined following Eq.
 (2), with , , and , respectively.
-The operating voltage for the central screening gate is handled separately as it has to satisfy two constraints: ensuring that all current channels can be formed without merging.
+The pinch-off occurs at roughly the inflection of the signal VP, the guaranteed isolation, VI, happens left of the pinch-off, and the operating voltages, VO, occur to the right of the pinch-off, typically on the slope.
+The operating voltage for the central screening gate is handled separately as it has to satisfy a secondary constraint: all current channels must be formed without merging.
 By default, .
-If at the subsequent state, a failure to form 1D current channels is observed twice, BATIS adjusts this parameter in increments defined as , where is chosen at random.
-The typical failure edge cases for this stage include too much depletion and too much accumulation; see the second row, the right column of Fig.
-2(c).
+If at the subsequent stage a failure to form 1D current channels is observed twice, BATIS adjusts this parameter in increments defined as , where is chosen at random.
+Typical unsatisfactory operational modes for this stage include too much depletion or accumulation; see the second row, the right column of Fig.
+3(c).
 If the channel below a gate is extraordinarily depleted, no voltage can be applied to the gate to accumulate it.
@@
 However, both types of issues would have been detected at the global accumulation stage, prompting the initiation of device illumination.
-Thus, the no change in current at this stage has to be caused either by an issue with the instrument or a broken gate.
-To check whether this is the case, the algorithm executes a repeated measurement.
+Thus, at this stage, the constant current signal is caused by either an instrument failure or a broken gate.
+To verify whether this is the case, BATIS executes a repeated measurement.
 If the error is observed a second time, it indicates an issue with the gate and the algorithm terminates.
-II.5 Forming 1D Current Channels
-The 1D current channels are formed by selectively removing 2DEG from certain regions of the device.
-It requires depleting the screening gates and accumulating the plungers and barriers (often collectively called finger gates) such that the current flows between the screening gates along the desirable path, as seen in the top row of Fig.
-2(d).
-Determining voltages where 1D channels are properly formed is typically done by sweeping two screening gates simultaneously in a 2D sweep while holding the finger gates constant.
+III.5 Current Channel Accumulation
+The 1D current channels are formed by selectively modifying the 2DEG in specific regions of the device.
+It involves depleting charges underneath the screening gates and accumulating them underneath the plunger and barrier gates (often collectively called finger gates) such that the current flows between the screening gates along the desired path, as seen in the top row of Fig.
+3(d).
+Determining voltages where 1D channels are properly formed is typically done iteratively, by sweeping two screening gates against each other in a 2D measurement while holding the corresponding finger gates constant.
 The formation of the channel manifests as a triangular-sloped region (the so-called triangle plots) [45].
-If the finger gates are not accumulated enough, this region will not form.
-Thus, multiple scans with incremented finger gate voltages—a three-dimensional dataset—are often required before the region is observed for all required channels.
-BATIS employs a different approach to simultaneously accumulate multiple 1D channels using only a single two-dimensional dataset.
-Instead of sweeping two screening gates, BATIS takes a 2D voltage scan while sweeping one screening gate against all of the finger gates between that gate and the other screening gate used to form the 1D channel (typically the central screening gate).
-BATIS uses the channel isolation (VI) and the operating points (VO) to prepare the device for characterizing the remaining screening gates by sweeping from VI past VP.
-The 1D channel formation technique implemented in BATIS requires a single measurement for each pair of gates, which makes it much more efficient than the standard approaches.
-The plot in the left panel in the second row Fig.
-2(d) exhibits three characteristic regions: a no-current region in the bottom left, a screening-gate-dominated region in the bottom right, and a finger-gate-dominated region in the top left.
-BATIS automatically selects an operating point in the top left part of the plot, near the boundary between the current and the no-current regions.
-The automatic analysis of the regions and their respective boundaries is enabled by fitting a synthetic model to the experimental data, similar to Ref.
-[32].
-Additional criteria based on the local properties, e.g., estimated resistance and derivatives of the measurement, are used to optimize the selection of the final point V for , among numerous candidates in the operating region.
+If the finger gates are not sufficiently accumulated, this region will fail to form.
+Multiple 2D scans with incremented finger gate voltages for each pair of screening gates—a three-dimensional dataset—are often necessary to observe the triangle region across all required channels.
+BATIS employs an alternative approach for initializing multiple 1D channels simultaneously.
+Rather than sweeping screening gates against each other, BATIS performs a 2D measurement of a screening gate against all the finger gates located between that gate and the opposing screening gate used to define the 1D channel (typically the central screening gate).
+Thus, this method efficiently captures the behavior of all channels in a single 2D dataset.
+A simulation of current as a function of gate voltages is shown in Fig.
+4(a), and in Fig.
+4(b) a corresponding example of experimental data is shown.
+Notably, there are four unique regions in the landscape in Fig.
+4(a): a no-current region in the bottom left, a finger-gate-dominated region in the top left (indicated with I), a screening-gate-dominated region in the bottom right (indicated with III), and a region of combined effect from both the fingers and screening gates (indicated with II).
+These regions all correspond to different operating modes of the device, with region I showing the proper accumulation of a 1D channel that is the required precursor to proper bootstrapping of the device; region II, in contrast, corresponds to an improper 2D channel formed under the screening and finger gates; and region III corresponds to an undesired 2D channel under the screening gate.
+All three of these regions are indicated in the simulation results shown in Fig.
+4(c).
 If the finger-gate-dominated region does not form, as shown in the right panel in the second row of Fig.
-2(d), BATIS illuminates the device with the finger gates voltages set to be more negative than the most recent illumination.
-This adjustment can be found in the configuration file, we used 400 .
-After such an illumination, the entire device response can change, because of cross-talk and the interdependence of the current paths.
-Thus, BATIS resumes autotuning from the global accumulation step.
-II.6 Finger Gate Characterization
-Once isolated current channels are formed, finger gate pinch-offs (VP) can be measured.
-Ideally, as the voltage on a given finger gate is lowered, the pinch-off curves should exhibit a smooth, sigmoidal trend consistent with Eq.
-1, eventually cutting off the current in the respective channels following Eq.
-(2) with .
-An idealized representation of such a curve is visualized in the left panel in the second row in Fig.
-2(e).
-The bottom panel in the third row of Fig.
-2(e) shows a typical curve for these measurements.
-As the voltage is reduced, the curve begins to exhibit the characteristic Coulomb blockade oscillations typical of QD devices [13].
-In principle, such oscillations could complicate any fit-based analysis.
-Instead, BATIS uses thresholding and slope detection to determine the pinch-off voltage on a point-by-point basis, going in the direction of increasing voltage (the reverse of how the signal was measured) to evade the small deviations of background noise, whilst capturing the first significant turn-on.
-III Experimental Demonstration
-We now demonstrate BATIS on a 1D quadruple Si/SixGe1-x QD device, nominally identical to the one shown in Fig.
-1(a) with 24 gates.
-The first layer of gates is made up of four screening gates, Si for , shown in Fig.
+3(d), BATIS initiates illumination with finger gates voltages set more negative than the most recent illumination.
+When the finger-gate-dominated region forms properly, BATIS automatically selects an operating point in region I, near the boundary between the current and the no-current regions.
+Fig.
+4(d) shows a Schrodinger-Poisson simulated example with properly set operating points for all current channels.
+The 1D channel formation technique implemented in BATIS requires a single measurement for each pair of screening/finger gates, which makes it much more efficient than the standard approaches.
+III.6 Finger Gate Characterization
+Once isolated current channels are formed, finger gate pinch-offs, VP, can be measured.
+Naively, one might hope that, as the voltage on a given finger gate is lowered, the pinch-off curves should exhibit a smooth, sigmoidal trend like that shown in the middle row of Fig.
+3(e).
+However, because finger gates are so small, trapped charge, e.g., in the oxide layer, often causes Coulomb blockade effects [26] to be visible, so that real devices display behavior like that shown in the third row of Fig.
+3(e).
+In this case, as the gate voltage is made less positive, the curve exhibits characteristic Coulomb blockade oscillations typical of QD devices.
+To be robust against this non-ideality, BATIS uses thresholding and slope detection during analysis to determine the pinch-off voltage on a point-by-point basis, going in the direction of increasing voltage from the end of the sweep (the reverse of how the signal was measured) to avoid small deviations of background noise, while capturing the first significant turn-on behavior.
+IV BATIS: Experimental Demonstration
+We demonstrate BATIS at 1.3 on a Si/SixGe1-x quadruple QD device with 24 gates, nominally identical to the one shown in Fig.
+1(a), which has two charge sensing channels in parallel with a channel to host the four qubit QDs.
+The first gate layer contains four screening gates, Si for , shown in Fig.
 1(a) in yellow.
 The next layer consists of five reservoirs, Ri for , and six plunger gates, Pi for .
-The reservoirs are shown in Fig.
-1(a) in blue, while the plungers are shown in red.
+The reservoirs are shown in blue in Fig.
+1(a), while the gates highlighted in red are the plunger gates.
 The final layer consists of nine barrier gates, Bi for , shown in Fig.
 1(a) in gray.
-The goal of BATIS is to form three current channels in the quad-QD device as it learned from the gate layout in the configuration, as indicated with white arrows in Fig.
-3(a).
-BATIS brings a grounded but not characterized device from Fig.
-1(a) to a fully operational device in Fig.
-1(d).
-In what follows, we explain how BATIS navigates the execution of the quad-QD device bootstrapping, testing, and characterizations.
+The goal of BATIS is to form three current channels and determine proper pinch-off gate voltages based on automated measurements and the gate layout information BATIS learns from the configuration file.
+This proper operation is indicated with black arrows in Fig.
+1(b), bringing a grounded but not-characterized device to a fully operational mode.
+In what follows, we explain how BATIS navigates the execution of the quad-QD device bootstrapping, testing, and characterization.
 First Leakage Test.
-BATIS begins by performing a device leakage test.
-For the quad-QD device used in our demonstration, the sample mount holding it contains connections, of which are used for instrument channels to control individual gates and ohmics and an additional wired channels are unused for this device.
-Thus, the size of the leakage matrix is to examine all possible leakage states.
+The sample mount holding the quad-QD device used here contains connections, of which are used for instrument channels to control individual gates and ohmics, and an additional wired channels are unused for this device.
+Thus, the size of the leakage matrix is to capture all possible leakage states.
+The leakage matrix is symmetric, and thus only the lower triangular part needs to be completed.
 BATIS first checks the diagonal elements of the leakage matrix.
-If the resistance of every connection exceeds 25, BATIS declares that there is no leakage in the device.
-If any of the diagonal elements falls below the 25 threshold, the device is declared leaking.
-Depending on where the current leak occurs, it is sometimes possible to recover a leaking device.
-Thus, rather than terminating, BATIS continues to fill in the leakage matrix for diagnostic purposes.
-Due to the symmetry of the measurements, the leakage matrix is also symmetric and only the lower triangular part of the matrix needs to be completed.
-Moreover, passing the threshold test by a diagonal component automatically guarantees that the entire column corresponding to that component also passes the test.
-Thus, only the columns corresponding to the elements below 25 need to be determined.
-BATIS continues to fill the leakage matrix one column at a time.
+If the resistance of any connection drops below 25 , BATIS declares that the device leaks.
+Moreover, if the threshold test fails on a diagonal component, BATIS automatically fills out the rest of the column for diagnostic purposes.
 In the worst case, a total of resistance measurements are necessary to complete the leakage matrix.
-The leakage test needs to be executed only once per cooldown if under fully automatic control.
-This is because BATIS has built-in safety checks during voltage choice selection to avoid exceeding safe voltage breakdowns between neighboring gates for all measurements.
-| Voltages | Run 1 | Run 2 | Run 3 | Run 4 | |
-| I1 | VP () | ||||
-| Vmax () | |||||
-| I2 | VP () | ||||
-| Vmax () | |||||
-| I3 | VP () | ||||
-| Vmax () |
+Under fully automatic control, the leakage test needs to be executed only once per cooldown.
+Damage following this test is unlikely, as BATIS incorporates built-in safety checks throughout its automated voltage selection process for all measurements to prevent dielectric breakdown between neighboring gates.
+| Voltages () | Run 1 | Run 2 | Run 3 | Run 4 | |
+| I1 | VP | ||||
+| Vmax | |||||
+| I2 | VP | ||||
+| Vmax | |||||
+| I3 | VP | ||||
+| Vmax |
 Global Accumulation.
-Next, BATIS performs the global accumulation stage.
-In our demonstration, the algorithm executed three rounds of illumination before the observed global turn-on values were determined to be satisfactory, occurring between and .
+Next, BATIS seeks to determine gate voltages that establish global accumulation.
+This step generates current flow across the device without regard to which channel or channels are connected to each other.
+In our demonstration, the algorithm executed three rounds of illumination before the observed global turn-on values were determined to be satisfactory, occurring in the desired range of to .
 By design, all gates are grounded during the first two illuminations.
-The first illumination is intended to completely reset the device while the second is to verify that all instruments work as intended.
-Because BATIS still did not find the gates in the desired range, it chose to apply voltages to all screening and reservoir gates for the third illumination.
-The voltages required to achieve turn-ons at around , calculated following the protocol in Ref.
-[21], were set to for gates associated with channel I1 and I2, and for gates associated with channel I3.
-The central screening gate, S2, was left grounded during the illumination.
-The fourth iteration of global turn-on measurement was successful.
+The first illumination is intended to completely reset the device, while the second is to verify that all instruments work as intended.
+Because BATIS still did not find the gates in the desired range for the third illumination, voltages were applied to all screening and reservoir gates.
+The voltages required to achieve turn-ons in the required range are determined by BATIS following the protocol in Ref.
+[34].
+In this case, during illumination, BATIS sets gates associated with channels I1 and I2 to , and gates associated with channel I3 to .
+The central screening gate, S2, was left grounded (i.e., set to 0 ) during illumination.
+Following the gate-biased illumination, the fourth iteration of the global turn-on measurement was successful.
 The summary of all four runs is presented in Table 1.
-The resulting configuration of the QD device is shown in the left panel in Fig.
-1(c), where the pinch-offs on all the screening gates, excluding S2, and all of the reservoir gates are colored based on the learned gate voltage.
+The resulting configuration of the QD device is shown in Fig.
+5(a), where the pinch-off voltages on all the screening gates, excluding S2, and all of the reservoir gates are colored based on the learned gate voltage.
 Accumulated Leakage Test.
@@
 Like previously, BATIS finds that all resistances exceed 25 .
-Since the ohmic connections are excluded, the resulting matrix has five fewer indexes.
-As none of the connections leaked, only measurements following the matrix diagonal were taken.
-Screening and Reservoir Characterization.
-The next stage in BATIS involves the characterization of the screening and reservoir gates.
+Since none of the connections leaked, only measurements following the matrix diagonal were taken.
+Screening and Reservoir Characterization The next step BATIS follows involves characterization of the screening and reservoir gates, which is important for dividing the induced 2DEG into three current channels.
 The results of these characterizations for all tested gates are presented in Fig.
-3(b).
-Unlike the other screening and reservoir gates, S2 needs to be measured simultaneously on channels I1, I2, and I3 since it borders all of the current channels.
-Because of this, the data for the S2 pinch-off can be found in Fig.
-3(c) as it needs to be organized by channel.
+6(b).
+The S2 gate borders all of the current channels and thus is not included here as it needs to be measured simultaneously on channels I1, I2, and I3.
+The data for the S2 pinch-off can be found in Fig.
+6(c).
+The voltage values are quite consistent between the sets of reservoir and screening gates intended to form the desired current channels, a feature that arises in part from the gate-biased illumination described above [34].
+We note that for gate S1, BATIS acquired the signal over a range, i.e., twice as long compared to all other screening gates.
+This is due to poor fit scoring using the reduced voltage scan range, which caused BATIS to perform additional measurements with an increased sweep range to ensure sufficient confidence in the derived pinch-off value.
+Since the long tail of the signal has no features, the signal shown in Fig.
+6(b) is truncated to be consistent with the remaining measurements for visualization purposes.
+The summary of the pinch-off voltages and the operating points selected by BATIS is presented in Table 2.
+A device in this configuration prepared for the channel accumulation stage is shown in Fig.
+5(b).
 | Gate | VP () | VI () | VO () |
@@
 | R5 |
-The voltage values are quite consistent between the sets of reservoir and screening gates intended to form the desired current channels, a feature from the illumination from earlier [21].
-We note that for gate S1, BATIS acquired the signal over a range, i.e., twice as long as compared to all other screening gates.
-This is likely due to poor fit scoring using the reduced voltage scan range, which caused BATIS to perform additional measurements with an increased sweep range to be more confident in the derived pinch-off value.
-Since the long tail of the signal has no features, the signal shown in Fig.
-3(b) is truncated to be consistent with the remaining measurements for visualization purposes.
-The summary of the pinch-off voltages and the operating points selected by BATIS is presented in Table 2.
-A device in this configuration prepared for the channel accumulation stage is shown in Fig.
-1(c).
-Forming 1D Current Channels.
-With all screening and reservoir gates characterized, BATIS proceeds to form the 1D current channels.
+Current Channel Accumulation With all screening and reservoir gates characterized and candidate operating voltages determined, BATIS proceeds to form the 1D current channels.
+As discussed above in Sec.
+III.5, the algorithm for 1D current channel formation implemented in BATIS does not follow the typical, triangle-plot-based strategy.
+Rather, it sweeps each of the three outer screening gates, i.e., S1, S3, and S4, against all finger gates relevant to the given 1D current channel.
+The traditional triangle plots are along a different set of axes (e.g., S2 and S1, instead of all finger gates and S1) than those shown in Fig.
+6(c).
+The advantage of measurements implemented in BATIS is faster convergence to the correct operating point for three well-formed and independent current channels.
+The automatic analysis of the regions and their respective boundaries is enabled by fitting a synthetic model to the experimental data, as explained in Ref.
+[46].
+Additional criteria based on the local properties, e.g., estimated resistance and derivatives of the measurement, are used to optimize the selection of the final operating point V for , among the numerous candidates in the operating region.
+If BATIS was unable to select an operating point after such a measurement, the next action would be to adjust all of the gates by using a 400 illumination.
+This adjustment can be found in the configuration file.
+After such an illumination, the entire device response can change because of crosstalk and the interdependence of the current paths.
+Thus, BATIS resumes autotuning from the global accumulation step.
 The quad-QD device requires three current channels to be formed, as indicated with the white arrows in Fig.
-3(a).
-BATIS identified three different operating points, one for each 1D current channel calculated from the simultaneous channel accumulation in Fig.
-3(c).
-The summary of the screening and finger gates configuration for each operating point is presented in Table 3, and a collapsed zero cross-talk visualization is pictured in the right panel in Fig.
-1(c).
-The three points are consolidated into one where the voltage color for each gate is the VO indexed by the channel the gate belongs to.
-As we discuss in Sec.II.5, the algorithm for 1D current channel formation implemented in BATIS does not follow the typical, triangle-plot-based strategy.
-Rather, it sweeps each of the three outer screening gates, i.e., S1, S3, and S4, against all finger gates relevant to the given 1D current channel.
-The traditional triangle plots [31, 32] are orthogonal to the measured planes shown in Fig.
-3(c) and would have been measurements with the S2 axis instead.
-In other words, the triangle plots and the plots utilized in BATIS contain similar information on different plane cuts through the same space.
+6(a).
+In this demonstration, BATIS identified three different operating points, one for each 1D current channel, based on the data shown in Fig.
+6(c).
+The summary of the screening and finger gates configuration for each operating point can be found in Table 3, and a collapsed visualization of these gate voltages is shown in Fig.
+5(c).
 | Gate(s) | V () | V () | V () |
@@
 The accumulation plots in Fig.
-3(c) reveal a dip in the current exiting when the current starts to flow in , shown by the dark green shape imposed on top of the signal in .
+6(c) reveal a dip in the current exiting when the current starts to flow in , shown by the dark green shape imposed on top of the signal in .
 This is due to a coupling between those two channels resulting from a shared ohmic connection to source the currents.
 Such a dip in the charge sensor readout is often called a shadow.
-Shadows increase the difficulty of detecting and forming isolated current channels.
-The accumulation plot for I1 is much less correlated with the current flow in the two charge sensor channels, but there is still some cross-talk.
-This is due to separation with the S2 gate which divides the 2DEG between the upper readout channels and the lower dot channel.
+Shadows increase the difficulty of detecting and forming isolated current channels, yet BATIS can handle this situation.
+In contrast, the accumulation plot for I1 is much less correlated with the current flow in the two charge sensor channels, but there is still some crosstalk arising from gate-to-2DEG cross capacitances.
+Finger Gate Characterization The last stage of BATIS involves the determination of pinch-off voltages for the finger gates.
+One of the problems with analyzing pinch-off data is the presence of spurious Coulomb blockade oscillations, as discussed above and shown in the bottom row in Fig.
+3(e).
+Such oscillations complicate fit-based analysis.
+Instead, BATIS uses thresholding and slope detection to determine the pinch-off voltage, searching in the increasing voltage direction (the reverse of how the signal was measured) to capture the first significant turn-on above the noise.
 | Gate | VP () | Gate | VP () | Gate | VP () |
@@
 | B5 | |||||
-Finger Gate Characterization.
-The last stage of BATIS involves finger gate characterization.
-This characterization is done individually for all of the finger gates on the QD device.
+This pinch-off voltage determination is performed individually for all of the finger gates of the QD device.
 The voltage range for each measurement is always the same and predefined in the configuration file.
-However, due to the Coulomb blockade oscillations, the automated search for the starting voltage for some gates is more challenging than for others.
-This can be seen in the raw data pictured in Fig.
-3(d).
+However, due to the Coulomb blockade oscillations, the automated search for the starting voltage for some gates is more challenging than for others, as is visible in the data shown in Fig.
+6(d), and the added complexity of the curves informed the algorithm described above and implemented here.
 The resulting pinch-off values of all of these varying curves for all finger gates are presented in Table 4.
 This configuration is used in the final colorized SEM for the ending state of the algorithm shown in Fig.
-1(d).
-IV Summary and outlook
-In this paper, we propose and demonstrate BATIS, a fully autonomous procedure for bootstrapping, automated testing, and characterization of accumulation mode undoped QD devices.
-The prior knowledge about the QD device required to initialize BATIS is limited to the heterostructure, the device geometry, the ohmic biases, and the device limits.
-BATIS navigates through complex high-dimensional gate space to find the final gate configuration in which all desirable current channels are formed following simple physical principles.
-The goal of BATIS is to test the device’s functionality to confirm that all gates can properly pinch-off and that all current channels can be formed.
-Testing and characterizing an unknown QD device is essential for identifying global device states that are suitable for observing the Coulomb blockade.
-Importantly, the single-button-press testing and initialization process is demonstrated at , eliminating the need for a full cooldown just to learn that a device is not functional.
-This result is particularly important as we envision a hot large-scale spin-based quantum processor.
-These systems will require qubit initialization, relying on a complete and accurate device characterization.
-One could imagine using BATIS to probe the attributes of many devices spread across a wafer or provide device diagnostics on a portion of a large qubit processor without going to temperatures.
-The flexible structure of BATIS allows for scalable characterization and initialization of any accumulation-mode linear array of QDs.
-It is also easily adaptable to other QD platforms.
-In particular, BATIS could straightforwardly mimic the behavior of algorithms proposed for nanowires and GaAs/AlxGa1-xAs by simply invoking skip logic on the steps not necessary for these devices.
-On the other hand, it is quite unlikely that any previously proposed algorithms could be easily adapted to the Si/SixGe1-x quad-QD device considered in this work.
+5(d).
+This voltage configuration describes the proper operating point in high-dimensional voltage space.
+BATIS has taken the device from an unknown initial state to a set of properly defined 1D channels with confirmed finger gate operation and a full compilation of pinch-off voltages.
+At this point, autotuners designed for 1D devices can be used on each individual channel, if desired [9].
+V Summary and outlook
+In this paper, we demonstrated an autonomous system for bootstrapping, automated testing, and characterization of gate-defined, accumulation-mode, undoped QD devices.
+This method, which we call BATIS, leverages knowledge from experiment and physical intuition—grounded in fundamental device physics—to electronically control navigation of the quantum device landscape.
+The prior knowledge about the QD device required to initialize BATIS is limited to the heterostructure type, the device geometry, the ohmic biases, and the device safety limits.
+The method proceeds in several steps, checking for device functionality as it moves forward.
+If irrecoverable device flaws are discovered, BATIS reports those issues and halts.
+Importantly, BATIS can recover from other non-idealities, such as undesirable amounts of trapped charge arising from fabrication or accident, by using illumination under gate-bias to modify in-situ the trapped charge and thus the device performance.
+This capability is important for resolving issues like charge disorder or incomplete turn-on, which may prevent functional behavior in otherwise operational devices.
+While much of the automated approach relies on existing experimental procedures, an innovation in BATIS is its method for 1D channel formation: rather than sweeping pairs of screening gates, BATIS performs a single 2D voltage scan between one screening gate and all intermediate finger gates.
+This strategy reduces the required measurement overhead to a single scan, resulting in substantial efficiency gains compared to conventional approaches.
+Importantly, the single-button-press testing and characterization process is demonstrated at , eliminating the need for a full cooldown for the device initialization.
+This result is particularly important as work advances towards “hot”, large-scale spin-based quantum processors.
+BATIS can also serve as a screening tool before cooling a dilution refrigerator to reach temperatures, if cold operation is desired.
 Previous studies have demonstrated device-agnostic algorithms for certain subsets of QD device-tuning tasks, such as setting QD topologies [47, 22] or charge-state classifications [14, 16].
-However, these algorithms presuppose that device characterization and foundational tuning steps—such as determining gate functionality, defining the QD channel(s), or locating regions of interest for the coarse-tuning regime—have already been performed.
-BATIS fills this gap, bringing device characterization easily within reach for linear QD arrays and paving the way toward fully automatic control.
-When linked together, the proposed subroutines form an autonomous system for controlling arbitrary QD devices up to any desired level of characterization, putting QD qubit systems on the path to scalable quantum computing.
+These algorithms presuppose that device characterization and foundational tuning steps—such as determining gate functionality, defining the QD channel(s), or locating regions of interest for the coarse-tuning regime—have already been performed.
+BATIS helps to fill this gap, enabling bootstrapping of gate-defined devices in which current paths are entirely defined by gate voltages whose values cannot be known in advance.
 Acknowledgements.
 We acknowledge Danielle Middlebrooks for helping with the automation.
-We acknowledge HRL Laboratories, LLC for support and L.F.
+We acknowledge HRL Laboratories, LLC, for support and L.F.
 Edge for providing one of the Si/SixGe1-x heterostructures used in this work.
@@
 Any mention of commercial products is for information only; it does not imply recommendation or endorsement by NIST.References
-- Burkard et al. [2023] G. Burkard, T. D. Ladd, A. Pan, J. M. Nichol, and J. R. Petta, Semiconductor spin qubits, Reviews of Modern Physics 95, 025003 (2023).
+- Burkard et al. [2023] G. Burkard, T. D. Ladd, A. Pan, J. M. Nichol, and J. R. Petta, Semiconductor spin qubits, Rev. Mod. Phys. 95, 025003 (2023).
 - van der Wiel et al.
@@
 R.
-Hulbert, et.
-al, A flexible design platform for si/sige exchange-only qubits with low disorder, Nano Lett.
-22, 1443 (2022), pMID: 34806894.
+Hulbert, et al., A flexible design platform for Si/SiGe exchange-only qubits with low disorder, Nano Lett.
+22, 1443 (2022) .
 - Wang et al.
@@
 L.
-Lawrie, et.
-al, Operating semiconductor quantum processors with hopping spins, Science 385, 447 (2024).
+Lawrie, et al., Operating semiconductor quantum processors with hopping spins, Science 385, 447 (2024).
 - George et al.
@@
 K.
-Harper, et.
-al, 12-spin-qubit arrays fabricated on a 300 mm semiconductor manufacturing line (2024), arXiv:2410.16583 [cond-mat.mes-hall] .
-- Note [2] For a given current curve, the pinch-off and the turn-on voltage are mathematically indistinguishable.
-However, when measuring a physical gate, a hysteresis (i.e., a shift of the current curve that depends on the direction of the sweep) can be observed.
-- Borsoi et al.
-[2024] F.
-Borsoi, N.
+Harper, et al., 12-spin-qubit arrays fabricated on a 300 mm semiconductor manufacturing line, arXiv:2410.16583 (2024) .
+- Zwolak and Taylor [2023] J. P. Zwolak and J. M. Taylor, Colloquium: Advances in automation of quantum dot devices control, Rev. Mod. Phys. 95, 011006 (2023).
+- Zwolak et al.
+[2024] J.
+P.
+Zwolak, J.
+M.
+Taylor, R.
 W.
-Hendrickx, V.
-John, M.
-Meyer, S.
-Motz, F.
-Van Riggelen, A.
-Sammak, S.
-L.
-De Snoo, G.
-Scappucci, and M.
-Veldhorst, Shared control of a 16 semiconductor quantum dot crossbar array, Nature Nanotechnology 19, 21 (2024).
-- Park et al.
-[2024] S.
-Park, J.
-Benson, J.
-Corrigan, J.
-P.
-Dodson, S.
-N.
-Coppersmith, M.
-Friesen, and M.
-A.
-Eriksson, Single-shot latched readout of a quantum dot qubit using barrier gate pulsing (2024), arXiv:2408.15380 [cond-mat].
-- Vandersypen et al.
-[2017] L.
-M.
-K.
-Vandersypen, H.
-Bluhm, J.
-S.
-Clarke, A.
-S.
-Dzurak, R.
-Ishihara, A.
-Morello, D.
-J.
-Reilly, L.
-R.
-Schreiber, and M.
-Veldhorst, Interfacing spin qubits in quantum dots and donors—hot, dense, and coherent, npj Quantum Inf.
-3, 34 (2017).
-- Meirav et al. [1990] U. Meirav, M. A. Kastner, and S. J. Wind, Single-electron charging and periodic conductance resonances in gaas nanostructures, Phys. Rev. Lett. 65, 771 (1990).
-- de Kruijf et al.
-[2023] M.
-de Kruijf, S.
-Geyer, T.
-Berger, M.
-Mergenthaler, F.
-Braakman, R.
-J.
-Warburton, and A.
-V.
-Kuhlmann, A compact and versatile cryogenic probe station for quantum device testing, Review of Scientific Instruments 94, 054707 (2023).
-- Neyens et al.
-[2024] S.
-Neyens, O.
-K.
-Zietz, T.
-F.
-Watson, F.
-Luthi, A.
-Nethwewala, H.
-C.
-George, E.
-Henry, M.
-Islam, A.
-J.
-Wagner, F.
-Borjans, et.
-al, Probing single electrons across 300-mm spin qubit wafers, Nature 629, 80 (2024).
-- Darulová et al.
-[2020] J.
-Darulová, S.
-J.
-Pauka, N.
-Wiebe, K.
+Andrews, J.
+Benson, G.
 W.
-Chan, G.
-C.
-Gardener, M.
-J.
-Manfra, M.
-C.
-Cassidy, and M.
-Troyer, Autonomous tuning and charge-state detection of gate-defined quantum dots, Phys.
-Rev.
-Appl.
-13, 054005 (2020).
-- Zubchenko et al.
-[2024] A.
-Zubchenko, D.
-Middlebrooks, T.
-Rasmussen, L.
-Lausen, F.
-Kuemmeth, A.
-Chatterjee, and J.
-P.
-Zwolak, Autonomous bootstrapping of quantum dot devices, arXiv:2407.20061 (2024).
-- Baart et al.
-[2016] T.
-A.
-Baart, P.
-T.
-Eendebak, C.
-Reichl, W.
-Wegscheider, and L.
-M.
-K.
-Vandersypen, Computer-automated tuning of semiconductor double quantum dots into the single-electron regime, Appl.
-Phys.
-Lett.
-108, 213104 (2016).
-- Lapointe-Major et al.
-[2020] M.
-Lapointe-Major, O.
-Germain, J.
-Camirand Lemyre, D.
-Lachance-Quirion, S.
-Rochette, F.
-Camirand Lemyre, and M.
-Pioro-Ladrière, Algorithm for automated tuning of a quantum dot into the single-electron regime, Phys.
-Rev.
-B 102, 085301 (2020).
-- Severin et al.
-[2024] B.
-Severin, D.
-T.
-Lennon, L.
-C.
-Camenzind, F.
-Vigneau, F.
-Fedele, D.
-Jirovec, A.
-Ballabio, D.
-Chrastina, G.
-Isella, M.
-De Kruijf, et.
-al, Cross-architecture tuning of silicon and SiGe-based quantum devices using machine learning, Sci.
-Rep.
-14, 17281 (2024).
-- Wolfe et al.
-[2024] M.
-A.
-Wolfe, B.
-X.
-Coe, J.
-S.
-Edwards, T.
-J.
-Kovach, T.
-McJunkin, B.
-Harpt, D.
-Savage, M.
-G.
-Lagally, R.
-McDermott, M.
-Friesen, et.
-al, Control of threshold voltages in Si/Si0.7 Ge0.3 quantum devices via optical illumination, Physical Review Applied 22, 034044 (2024).
+Bryant, D.
+Buterakos, A.
+Chatterjee, S.
+Das Sarma, M.
+Eriksson, E.
+Greplová, et al., Data needs and challenges for quantum dot devices automation, npj Quantum Inf.
+10, 105 (2024).
+- Note [2] Mathematically, the pinch-off and the turn-on voltage are indistinguishable for a given current curve. However, when measuring a physical gate, a hysteresis can be observed.
 - Zwolak et al.
@@
 Taylor, Ray-based framework for state identification in quantum dot devices, PRX Quantum 2, 020335 (2021).
+- Durrer et al.
+[2020] R.
+Durrer, B.
+Kratochwil, J.
+Koski, A.
+Landig, C.
+Reichl, W.
+Wegscheider, T.
+Ihn, and E.
+Greplova, Automated tuning of double quantum dots into specific charge states using neural networks, Phys.
+Rev.
+Appl.
+13, 054019 (2020).
+- Lapointe-Major et al.
+[2020] M.
+Lapointe-Major, O.
+Germain, J.
+Camirand Lemyre, D.
+Lachance-Quirion, S.
+Rochette, F.
+Camirand Lemyre, and M.
+Pioro-Ladrière, Algorithm for automated tuning of a quantum dot into the single-electron regime, Phys.
+Rev.
+B 102, 085301 (2020).
 - Ziegler et al.
@@
 P.
-Zwolak, MAViS: Modular Autonomous Virtualization System for Two-Dimensional Semiconductor Quantum Dot Arrays, arXiv:2411.12516 (2024).
-- Zwolak and Taylor [2023] J. P. Zwolak and J. M. Taylor, Colloquium: Advances in automation of quantum dot devices control, Rev. Mod. Phys. 95, 011006 (2023).
+Zwolak, Modular Autonomous Virtualization System for Two-Dimensional Semiconductor Quantum Dot Array, Phys.
+Rev.
+X 15, 021034 (2025).
+- Baart et al.
+[2016] T.
+A.
+Baart, P.
+T.
+Eendebak, C.
+Reichl, W.
+Wegscheider, and L.
+M.
+K.
+Vandersypen, Computer-automated tuning of semiconductor double quantum dots into the single-electron regime, Appl.
+Phys.
+Lett.
+108, 213104 (2016).
+- Darulová et al.
+[2020] J.
+Darulová, S.
+J.
+Pauka, N.
+Wiebe, K.
+W.
+Chan, G.
+C.
+Gardener, M.
+J.
+Manfra, M.
+C.
+Cassidy, and M.
+Troyer, Autonomous tuning and charge-state detection of gate-defined quantum dots, Phys.
+Rev.
+Appl.
+13, 054005 (2020).
+- Severin et al.
+[2024] B.
+Severin, D.
+T.
+Lennon, L.
+C.
+Camenzind, F.
+Vigneau, F.
+Fedele, D.
+Jirovec, A.
+Ballabio, D.
+Chrastina, G.
+Isella, M.
+De Kruijf, et al., Cross-architecture tuning of silicon and SiGe-based quantum devices using machine learning, Sci.
+Rep.
+14, 17281 (2024).
+- Zubchenko et al.
+[2024] A.
+Zubchenko, D.
+Middlebrooks, T.
+Rasmussen, L.
+Lausen, F.
+Kuemmeth, A.
+Chatterjee, and J.
+P.
+Zwolak, Autonomous bootstrapping of quantum dot devices, Phys.
+Rev.
+Appl.
+23, 014072 (2025).
+- Borsoi et al.
+[2023] F.
+Borsoi, N.
+W.
+Hendrickx, V.
+John, M.
+Meyer, S.
+Motz, F.
+van Riggelen, A.
+Sammak, S.
+L.
+de Snoo, G.
+Scappucci, and M.
+Veldhorst, Shared control of a 16 semiconductor quantum dot crossbar array, Nat.
+Nanotechnol.
+19, 21 (2023).
+- Park et al.
+[2024] S.
+Park, J.
+Benson, J.
+Corrigan, J.
+P.
+Dodson, S.
+N.
+Coppersmith, M.
+Friesen, and M.
+A.
+Eriksson, Single-shot latched readout of a quantum dot qubit using barrier gate pulsing, arXiv:2408.15380 (2024).
+- Vandersypen et al.
+[2017] L.
+M.
+K.
+Vandersypen, H.
+Bluhm, J.
+S.
+Clarke, A.
+S.
+Dzurak, R.
+Ishihara, A.
+Morello, D.
+J.
+Reilly, L.
+R.
+Schreiber, and M.
+Veldhorst, Interfacing spin qubits in quantum dots and donors—hot, dense, and coherent, npj Quantum Inf.
+3, 34 (2017).
+- Meirav et al. [1990] U. Meirav, M. A. Kastner, and S. J. Wind, Single-electron charging and periodic conductance resonances in GaAs nanostructures, Phys. Rev. Lett. 65, 771 (1990).
+- de Kruijf et al.
+[2023] M.
+de Kruijf, S.
+Geyer, T.
+Berger, M.
+Mergenthaler, F.
+Braakman, R.
+J.
+Warburton, and A.
+V.
+Kuhlmann, A compact and versatile cryogenic probe station for quantum device testing, Rev.
+Sci.
+Instrum.
+94, 054707 (2023).
+- Neyens et al.
+[2024] S.
+Neyens, O.
+K.
+Zietz, T.
+F.
+Watson, F.
+Luthi, A.
+Nethwewala, H.
+C.
+George, E.
+Henry, M.
+Islam, A.
+J.
+Wagner, F.
+Borjans, et al., Probing single electrons across 300-mm spin qubit wafers, Nature 629, 80 (2024).
+- Paurevic et al.
+[2025] A.
+Paurevic, A.
+Sakr, T.
+Joshi, D.
+van der Bovenkamp, Q.
+T.
+Nicolau, F.
+A.
+Zwanenburg, and J.
+Baugh, Automated tuning and characterization of a single-electron transistor charge sensor, arXiv:2502.02521 (2025) .
+- Sze and Ng [2006] S. M. Sze and K. K. Ng, Physics of Semiconductor Devices (Wiley, New York, 2006).
+- Massai et al.
+[2024] L.
+Massai, B.
+Hetényi, M.
+Mergenthaler, F.
+J.
+Schupp, L.
+Sommer, S.
+Paredes, S.
+W.
+Bedell, P.
+Harvey-Collard, G.
+Salis, A.
+Fuhrer, and N.
+W.
+Hendrickx, Impact of interface traps on charge noise and low-density transport properties in Ge/SiGe heterostructures, Commun.
+Mater.
+5, 151 (2024).
+- Rudolph et al. [2019] M. Rudolph, B. Sarabi, R. Murray, M. S. Carroll, and N. M. Zimmerman, Long-term drift of Si-MOS quantum dots with intentional donor implants, Sci. Rep. 9, 7656 (2019).
+- Meyer et al.
+[2023] M.
+Meyer, C.
+Déprez, I.
+N.
+Meijer, F.
+K.
+Unseld, S.
+Karwal, A.
+Sammak, G.
+Scappucci, L.
+M.
+K.
+Vandersypen, and M.
+Veldhorst, Single-electron occupation in quantum dot arrays at selectable plunger gate voltage, Nano Lett.
+23, 11593 (2023) .
+- Wolfe et al.
+[2024a] M.
+Wolfe, B.
+X.
+Coe, J.
+S.
+Edwards, T.
+J.
+Kovach, T.
+McJunkin, B.
+Harpt, D.
+Savage, M.
+Lagally, R.
+McDermott, M.
+Friesen, et al., Control of threshold voltages in /0.70.3 quantum devices via optical illumination, Phys.
+Rev.
+Appl.
+22, 034044 (2024a).
 - Zajac et al.
@@
 Eriksson, Fabrication process and failure analysis for robust quantum dots in silicon, Nanotechnology 31, 505001 (2020).
+- [37] See Supplemental Material at URL-will-be-inserted-by-publisher.
+- Wang et al.
+[2022] K.
+Wang, G.
+Xu, F.
+Gao, H.
+Liu, R.-L.
+Ma, X.
+Zhang, Z.
+Wang, G.
+Cao, T.
+Wang, J.-J.
+Zhang, et al., Ultrafast coherent control of a hole spin qubit in a germanium quantum dot, Nat.
+Commun.
+13, 206 (2022).
+- Tormo-Queralt et al.
+[2022] R.
+Tormo-Queralt, C.
+B.
+Møller, D.
+A.
+Czaplewski, G.
+Gruber, M.
+Cagetti, S.
+Forstner, N.
+Urgell-Ollé, J.
+A.
+Sanchez-Naranjo, C.
+Samanta, C.
+S.
+Miller, and A.
+Bachtold, Novel nanotube multiquantum dot devices, Nano Lett.
+22, 8541 (2022) .
+- Dirac and Fowler [1926] P. A. M. Dirac and R. H. Fowler, On the theory of quantum mechanics, Proc. R. Soc. Lond. Ser. A-Contain. Pap. Math. Phys. Character. 112, 661 (1926).
+- Zannoni [1999] A. Zannoni, On the quantization of the monoatomic ideal gas, arXiv:cond-mat/9912229 (1999).
+- Bardeen [1947] J. Bardeen, Surface states and rectification at a metal semi-conductor contact, Phys. Rev. 71, 717 (1947).
+- Dacey and Ross [1953] G. C. Dacey and I. M. Ross, Unipolar "field-effect" transistor, Proc. IRE 41, 970 (1953).
 - Note [3] The parameter is defined as part of the configuration file and can be adjusted if needed.
@@
 P.
-Zwolak, Automation of Quantum Dot Measurement Analysis via Explainable Machine Learning (2024), version Number: 4.
+Zwolak, Automation of quantum dot measurement analysis via explainable machine learning, Mach.
+Learn.: Sci.
+Technol.
+6, 015006 (2025).
 - Ziegler et al.
@@
 17, 024069 (2022).
-- Durrer et al.
-[2020] R.
-Durrer, B.
-Kratochwil, J.
-Koski, A.
-Landig, C.
-Reichl, W.
-Wegscheider, T.
-Ihn, and E.
-Greplova, Automated tuning of double quantum dots into specific charge states using neural networks, Phys.
-Rev.
-Appl.
-13, 054019 (2020).
+See pages ,1,,2,,3 of supp.pdf